### v1

# Predicting Multi-Agent Specialization via Task Parallelizability

Abstract
Multi-agent systems often rely on specialized agents with distinct roles rather than general-purpose agents that perform the entire task independently. However, the conditions that govern the optimal degree of specialization remain poorly understood. In this work, we propose that specialist teams outperform generalist ones when environmental constraints limit task parallelizability—the potential to execute task components concurrently. Drawing inspiration from distributed systems, we introduce a heuristic to predict the relative efficiency of generalist versus specialist teams by estimating the speed-up achieved when two agents perform a task in parallel rather than focus on complementary subtasks. We validate this heuristic through three multi-agent reinforcement learning (MARL) experiments in Overcooked-AI, demonstrating that key factors limiting task parallelizability influence specialization. We also observe that as the state space expands, agents tend to converge on specialist strategies, even when generalist ones are theoretically more efficient, highlighting potential biases in MARL training algorithms. Our findings provide a principled framework for interpreting specialization given the task and environment, and introduce a novel benchmark for evaluating whether MARL finds optimal strategies.
multi-agent reinforcement learning, collaboration, efficiency, specialization, policy diversity
1 Introduction
Collaboration enables multi-agent systems to achieve goals that exceed the capabilities of a single agent. Advances in artificial intelligence (AI) have enhanced the capabilities of multi-agent systems to complete complex, coordinated tasks, including automating warehouses (Canese et al., 2021), sustainable resource management (Perolat et al., 2017), and simulated social interactions (Park et al., 2023). However, efficient task allocation remains a key challenge.
Existing work across multiple fields, including multi-agent reinforcement learning (MARL), cognitive science, and robotics, often assumes that specialization is universally desirable (Padgham & Winikoff, 2002; Zhu & Zhou, 2008; Charakorn et al., 2020; Wang et al., 2020a, b; Li et al., 2021; Bettini et al., 2023; Juang et al., 2024; Bettini et al., 2024; Swanson et al., 2024). For example, fine dining restaurants rely on specialists—the saucier prepares sauces, the rôtisseur roasts, and so on—who each perform distinct yet complementary subtasks. Intuitively, specialization narrows action spaces (Wang et al., 2020b), reduces ambiguity about what agents should do next (Goldstone et al., 2024), and eases individual cognitive loads (Griffiths, 2020). This frames task allocation as an optimization problem, focusing on identifying roles and assigning agents accordingly (Wang et al., 2020a).
However, human teams demonstrate that specialization is not always optimal. Fine dining chefs coordinate to produce complex meals, but at the cost of high training overhead, rigidity, and failure risk if a key specialist is unavailable (Ben-Oren et al., 2023; Vélez et al., 2024). In contrast, generalist fry cooks perform the same subtasks in parallel, such as grilling and assembling, maximizing adaptability and throughput. Examples of specialist and generalist strategies are depicted in Figure 1. Furthermore, although AI agents also face coordination challenges (Shoham & Leyton-Brown, 2008; Wu et al., 2021), it remains unclear whether the constraints that make specialization advantageous for people also apply. With agents becoming increasingly capable we must revisit this assumption and ask: under what conditions, if any, is specialization truly optimal?
Figure 1: Specialists vs. generalists. (A) Specialist teams tend to focus on non-overlapping subtasks. In Overcooked, an example of this is Agent 1 putting onions and tomatoes in the pot, while Agent 2 fetches a bowl and serves the soup. In highly specialized teams, agents’ high-level action distributions are completely distinct, resulting in a Jensen-Shannon divergence ($\text{JSD}(P_{1}||P_{2})$) of 1.0. Example specialist video. (B) In generalist teams, both agents perform all subtasks. In Overcooked, this may look like Agents 1 and 2 independently making their own soups in parallel. In generalist teams, agents’ action distributions are the same, resulting in $\text{JSD}(P_{1}||P_{2})=0.0$. Example generalist video.
We propose that a task’s parallelizability—the potential for its independent subtasks to be executed simultaneously—limits the benefits of generalist teams. Drawing from the distributed computing literature (Almasi & Gottlieb, 1994; McCool et al., 2012; Vélez et al., 2023), we use Amdahl’s Law to determine when multi-agent systems should favor generalists or specialists. Amdahl’s Law predicts that there are diminishing returns to increasing the number of processors performing a task in parallel, based on the portion of the task that must be performed serially (Amdahl, 1967; Hill & Marty, 2008; Cassidy & Andreou, 2011). Applied to multi-agent systems, this places an upper bound on performance gains achievable from generalist agents working in parallel. When spatial and resource constraints limit parallel execution, agents should improve efficiency by specializing in complementary subtasks. Using a simple, interpretable model, we provide a novel framework for evaluating the trade-offs between specialist and generalist strategies.
To test this prediction, we develop a theoretical framework for comparing generalist and specialist efficiency and apply it to three multi-agent experiments in Overcooked-AI. Experiment 1 shows that key model variables influence specialization, with a moderate negative correlation between predicted task parallelizability and specialization, though larger state spaces complicate this relationship. Experiment 2 controls for these confounds, demonstrating our model’s high accuracy in predicting specialization. Experiment 3 finds that larger state spaces promote specialization by increasing coordination challenges. Our findings suggest that specialization is driven by task parallelizability and environment, enabling the design of layouts that encourage or inhibit policy diversity. More broadly, our framework highlights the value of using insights from distributed computing to improve the interpretability of multi-agent systems.
2 Theoretical Framework
We introduce a general framework for evaluating task parallelizability and predicting emergent specialization in multi-agent systems.
2.1 $N$-Player Markov Game
Following Carroll et al. (2019) and Ruhdorfer et al. (2024), we formalize a cooperative decentralized Markov decision process (MDP), defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{N},A,\mathcal{T},\mathcal{R},\gamma\rangle$. $\mathcal{S}$ is a finite set of states, and $\mathcal{N}$ is a set of $N$ agents. $A$ is the joint action space, where each agent $i\in\mathcal{N}$ selects an action $a_{i}\in A$ to form a joint action $a=(a_{1},a_{2},...,a_{N})$. $\mathcal{T}:\mathcal{S}\times A\to\Delta(\mathcal{S})$ is the transition function which defines the probability distribution over next states given the current state and joint action. Finally, $\mathcal{R}:\mathcal{S}\times A\to\mathbb{R}$ is the shared reward function. The objective of each agent is to maximize its expected discounted return $\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\big{]}$, where $\gamma\in[0,1)$ is a discount factor and $r(s_{t},a_{t})$ is the shared reward received by all agents at time $t$. While agents share the same reward, each optimizes its own cumulative return independently.
2.2 Jensen-Shannon Divergence and Specialization
We define specialization as the differentiation of agents’ policies, capabilities, or roles in a multi-agent task. Let $\pi^{i}:\mathcal{S}\to\Delta(\mathcal{A}^{i})$ be the policy of agent $i$ that maps each state $s\in\mathcal{S}$ to a distribution over actions $a\in\mathcal{A}^{i}$. In generalist teams, agents do not specialize in specific subtasks, and thus select from the entire set of actions $A^{i}$. Formally, $\forall a\in\mathcal{A}^{i},\pi_{generalist}^{i}(a|s)>0\text{ for some }s\in\mathcal{S}$. In specialist teams, agents are restricted to specific subtasks and resources. In this case, $\mathcal{A}_{specialist}^{i}\subset\mathcal{A}^{i}$, and their specialist policies satisfy $\forall a\notin\mathcal{A}_{specialist}^{i},\pi_{specialist}^{i}(a|s)=0$. Examples of action distributions and specialist versus generalist strategies can be found in Figure 1.
Let the set of high-level actions be denoted as $\mathcal{A}=\{a_{1},a_{2},...,a_{k}\}$ (Amato et al., 2019). Each agent follows a state-visitation distribution that describes how frequently they occupy different states under their policy (Schulman et al., 2015) given by:
$d^{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s|\pi)$ (1)
The corresponding state-action visitation distribution is:
$d^{\pi}(s,a)=d^{\pi}(s)\pi(a|s),$ (2)
which represents the frequency with which an agent selects action $a$ in state $s$ under policy $\pi$. For each agent $i$, we define their effective action distribution using their visitation distribution:
$P_{i}(a)=\sum_{s\in\mathcal{S}}d^{\pi}(s,a)$ (3)
This distribution captures both what actions an agent tends to take and how often they visit states where those actions are available. We approximate specialization using the Jensen-Shannon divergence (JSD) over agents’ action distributions $P_{i}$ (Endres & Schindelin, 2003; Fuglede & Topsoe, 2004). Given $N$ agents with action distributions $P_{1},P_{2},...,P_{N}$, JSD is given by:
$\text{JSD}(P_{1},P_{2},...,P_{N})=\sum_{i=1}^{N}w_{i}D_{\mathrm{KL}}(P_{i}||M),$ (4)
where $M=\sum_{i=1}^{N}w_{i}P_{i}$ is the mixture distribution, $w_{i}=\frac{1}{N}$, and $D_{\mathrm{KL}}(P_{i}||M)$ is the Kullback-Leibler (KL) divergence between each agent’s action distribution $P_{i}$ and $M$. When $\text{JSD}(P_{1},P_{2},...,P_{N})=0.0$, all $N$ agents have performed the same proportions of actions, indicating that all agents have performed all subtasks equally (no specialization). When $\text{JSD}(P_{1},P_{2},...,P_{N})=1.0$, the action distributions are completely disjoint (i.e., each agent assigns probability exclusively to a unique set of actions with no overlap), meaning that the agents specialize in complementary subtasks. Intermediate values of JSD indicate partial specialization, in which the agents overlap in some subtasks but not others. We use Jensen-Shannon rather than Kullback-Leibler divergence because it is symmetric, i.e. $\text{JSD}(P_{1}||P_{2})=\text{JSD}(P_{2}||P_{1})$. The degree of specialization in each team should be the same regardless of which agent specializes in which subtasks.
2.3 Graphical Representations of Spatial Layouts
To model task execution in structured environments, we represent spatial layouts as graphs. This allows for efficient computation of movement paths, bottlenecks, and spatial constraints on multi-agent coordination. We define a layout graph $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of walkable positions and task-relevant objects and $\mathcal{E}$ is the set of valid movements between adjacent positions. This spatial representation applies broadly to spatio-temporal multi-agent environments, including gridworlds. $G$ defines the static structure of the environment and constrains the MDP, where the state space $\mathcal{S}$ includes agent positions $x\subseteq\mathcal{V}$ and task-relevant object states. The transition function $\mathcal{T}$ ensures that agents execute valid movements along edges in $\mathcal{E}$.
Definition 2.1.
A layout graph is an unweighted, directed graph $G=(\mathcal{V},\mathcal{E})$ where:
$V=\{u\mid u\in\{(i,j)\mid i\in[1,n],j\in[1,m]\}\setminus\text{obstacles}\}$
An edge $(u,v)\in\mathcal{E}$ exists if and only if $u$ and $v$ are adjacent, i.e., they share a horizontal or vertical edge in the grid. The type of edge between $u$ and $v$ depends on the type of nodes: (i) If $u$ and $v$ are both walkable nodes, there is a bidirectional edge such that $(u,v)\in\mathcal{E}$ and $(v,u)\in\mathcal{E}$. (ii) If $v$ is a task-relevant object (e.g., workstation) there is a unidirectional edge $(u,v)\in\mathcal{E}$, but $(v,u)\notin\mathcal{E}$, reflecting the fact that agents can interact with but not move through it.
2.4 Task Decomposition as a Directed Acyclic Graph
To analyze task parallelizability, we represent task execution as a directed acyclic graph (DAG), where nodes denote subtasks and edges represent the order in which subtasks must be completed (Solway et al., 2014).
Definition 2.2.
We define a task graph $T=\{\mathcal{B}_{r},\mathcal{E}_{T}\}$, where $\mathcal{B}_{r}$ (subtask nodes) represent intermediate steps such as workstations or resources, and $\mathcal{E}_{T}$ indicate ordering constraints between subtasks (e.g., an agent must retrieve an item before assembling it).
The number of subtasks $m$ depends on task complexity. Using the layout graph $G$, we determine the spatial locations and paths between subtasks. The fraction $f(i)$ of the total task spent on a given subtask $i$ is:
$f(i)=\frac{d(i)}{\sum_{j=1}^{m}d(j)},$ (5)
where $d(i)$ is the path length required to complete subtask $i$.
2.5 Spatial and Resource Bottlenecks
We introduce a framework to identify bottlenecks that constrain parallelizability in multi-agent tasks. This idea is inspired by sequential dependencies, resource contention, and communication overhead that prevent scalability in parallel processing (Almasi & Gottlieb, 1994; McCool et al., 2012; Zhuravlev et al., 2010). A classic example is memory bandwidth in GPUs, where increasing the number of threads does not lead to proportional speed-up due to limited shared access (Dublish et al., 2017). We extend this framework to multi-agent tasks by identifying two key types of bottlenecks: (1) spatial bottlenecks ($\mathcal{B}_{s}$) limit movement of multiple agents due to obstacles, narrow passages, and congestion, and (2) resource bottlenecks ($\mathcal{B}_{r}$) limit access to critical resources such as workstations or tools needed to complete subtasks (Mieczkowski et al., 2024). Together, these bottlenecks determine how effectively agents can execute tasks in parallel. They arise across numerous multi-agent domains, including warehouse operations (insufficient space, storage locations, and pallets (Živičnjak et al., 2022)), autonomous vehicles (lane reductions, traffic congestion (Zhang & Gao, 2020)), and smart factories (Wang et al., 2016).
We represent spatial and resource bottlenecks as two global sets over $G=(\mathcal{V},\mathcal{E})$: $\mathcal{B}_{s}\subseteq\mathcal{E}$, $\mathcal{B}_{r}\subseteq\mathcal{V}$, where $\mathcal{B}_{s}$ consists of edges in $\mathcal{E}$ that constrain movement and $\mathcal{B}_{r}$ consists of nodes in $\mathcal{V}$ with task-relevant objects (e.g., workstations). Each subtask $i$ is constrained by a subset of bottlenecks relevant to the path ($\mathcal{B}_{s}(i)$) and resources ($\mathcal{B}_{r}(i)$) required to complete it.
To quantify spatial bottlenecks, we compute edge betweenness centrality, a graph-theoretic measure of congestion where high-centrality edges indicate critical bottlenecks in connectivity (Barrat et al., 2004).
Definition 2.3.
The betweenness centrality of an edge $e$ in graph $G=(\mathcal{V},\mathcal{E})$ is:
$C_{B}(e)=\sum_{s\neq t\in V}\frac{\phi(s,t|e)}{\phi(s,t)},$ (6)
where $\phi(s,t)$ is the number of shortest paths between nodes $s$ and $t$, and $\phi(s,t|e)$ is the number of those paths that pass through edge $e$.
Since not all bottlenecks affect every subtask, we compute a per-subtask spatial bottleneck score by summing the centralities of all edges along the shortest path required for that subtask: $\mathcal{B}_{s}(i)=\sum_{e\in\text{path}(u,v)}C_{B}(e)$, where path$(u,v)$ is the shortest path between subtask-relevant nodes $u$ and $v$. A higher $\mathcal{B}_{s}(i)$ indicates stronger spatial constraints that may inhibit parallel agent movement.
To quantify resource bottlenecks, we simply sum the number of task-relevant nodes that are required for a subtask $i$. For example, in a manufacturing system there may be only two welding stations; thus, only two agents may complete that subtask at the same time. Formally, for each subtask $i$, we define a resource bottleneck score $\mathcal{B}_{r}(i)$ as: $\mathcal{B}_{r}(i)=\sum_{v\in\mathcal{B}_{r(i)}}c(v)$, where $c(v)$ is the capacity of each task-relevant node (or the maximum agents that can use it at once).
2.6 Task Parallelizability Prediction
We now adapt Amdahl’s Law, a measure of the expected speed-up when dividing a task across multiple processors given the proportion that must be executed serially, to multi-agent tasks (Amdahl, 1967; Hill & Marty, 2008; Cassidy & Andreou, 2011). For the original form of Amdahl’s Law, see Appendix A.1. Mieczkowski et al. (2024) extended this idea to predict idleness during human collaborations given the current group size, workload, and resource constraints. We hypothesize that a similar formulation can be used to predict specialization in multi-agent tasks.
Let $f(i)$ represent the fraction of the total task time required for the $i^{th}$ of $m$ subtasks such that $\sum_{i=1}^{m}f(i)=1$. For each subtask $i$, the overall bottleneck capacity is defined as the minimum bottleneck that prevents agents executing it concurrently, or $\mathcal{B}_{i}=\min(\mathcal{B}_{s}(i),\mathcal{B}_{r}(i))$. The subtask speed-up factor $s_{i}(N,\mathcal{B}_{i})$ quantifies the efficiency gain from adding an $N$th agent to a subtask, given that $N-1$ agents are already working on it and $\mathcal{B}_{i}$ bottlenecks restrict parallel execution. Formally:
$s(N,\mathcal{B}_{i})=\frac{\text{min}(N,\mathcal{B}_{i})}{\text{min}(N-1,\mathcal{B}_{i})}$ (7)
Thus, if we assume that there is no cost to switching between subtasks, the overall expected speed-up achievable when an additional agent performs the entire task in parallel, which we call task parallelizability, is expressed as:
$S(N,\mathcal{B})=\frac{1}{\sum_{i=1}^{m}\frac{f(i)}{s(N,\mathcal{B}_{i})}}$ (8)
2.7 Task Parallelizability and Specialization
Our task parallelizability prediction estimates the efficiency gain if an additional agent independently completed all $m$ subtasks in parallel. We propose that this measure sets a fundamental upper bound on the benefits of specialization in multi-agent systems. If $S(N,\mathcal{B})$ is low due to resource or spatial bottlenecks, then the task would not be significantly sped up by multiple agents executing it in parallel. In this case, there would be little value to a team of generalist agents, so if we want to increase task efficiency it is better for agents to specialize and perform complementary subtasks. However, specialization itself can introduce new bottlenecks, particularly when agents must coordinate frequently or wait for specific subtasks to be completed before proceeding. In environments with high parallelizability $S(N,\mathcal{B})$, generalist agents who work independently may avoid these delays, leading to faster performance.
3 Application to Overcooked-AI
To investigate how task parallelizability influences specialization, we apply our theoretical model for predicting generalist efficiency to a specific multi-agent environment (Carroll et al., 2019). Overcooked-AI is a fully cooperative two-agent platform in which pairs of agents must deliver soups as quickly as possible while navigating obstacles, avoiding collisions, and coordinating subtasks (Wu et al., 2021). Each recipe requires a different number of subtasks ($m$) to prepare a soup, while resource bottlenecks such as the number of pots ($\mathcal{B}_{r}$) and spatial bottlenecks such as narrow passages ($\mathcal{B}_{s}$) constrain task parallelizability.
Layout Graphs. We represent each layout graph according to Definition 2.1. An example is depicted in Figure 2.
Figure 2: Examples of Overcooked layouts, layout graphs, and spatial bottleneck scores $\mathcal{B}_{s}(i)$ for subtask Onion $\to$ Pot. Layout A is an open room with low spatial bottlenecks, whereas B has a narrow passage leading to the pot, creating a greater bottleneck.
Task Decomposition. We decompose the overall task of “cooking a soup” using a directed acyclic graph (DAG) $T=(\mathcal{B}_{r},\mathcal{E}_{T})$ according to Definition 2.2, where $\mathcal{B}_{r}$ is the set of resources, or workstations (e.g., onion piles, pots, bowls), and $(m_{i}\to m_{j})\in\mathcal{E}_{T}$ is a directed edge representing a required subtask (such as moving an onion to the pot). An example can be found in Figure 5. The number of subtasks $m$ depends on the recipe. To make a one-onion soup, there are $m=4$ subtasks total: Onion $\to$ Pot, Cook in Pot, Bowl $\to$ Pot, and Pot $\to$ Serve. To make a two-onion or one-onion one-tomato soup, there are $m=5$ subtasks. To make a three-onion or two-onion one-tomato soup, there are $m=6$ subtasks. Using the layout graph $G$, we determine the locations and paths between workstations. The fraction $f(i)$ spent on subtask $i$ is its path length relative to the total path length of all subtasks following Equation 5.
Bottlenecks. We manipulate and measure bottlenecks in Overcoooked in two ways. We vary resource bottlenecks $\mathcal{B}_{r}$ by changing the number of workstations ($p$ pots, $o$ onions, $t$ tomatoes, $d$ bowls, and $s$ serving stations) available to agents. We vary spatial bottlenecks $\mathcal{B}_{s}$ by modifying the layout size and arrangement of obstacles (represented as counters that agents cannot cross). We quantify these spatial bottlenecks for each subtask $i$ using edge betweenness centrality as described in Definition 2.3.
Specialization. To measure specialization, we convert agents’ behaviors (move, interact, wait) from a policy roll-out of the best-performing seed to high-level actions: (i) picking up onions, (ii) placing onions in pot, (iii) picking up tomatoes, (iv) placing tomatoes in pot, (v) picking up bowls, (vi) moving soup from pot to bowl, and (vii) serving soup. Each agent $i\in\mathcal{N}$ has an action distribution $P_{i}:\mathcal{A}\to[0,1]$, where $P_{i}(a)$ represents the probability of agent $i$ taking action $a\in\mathcal{A}$ during a policy roll-out after training. For each agent $i$, we compute their high-level action distribution $P_{i}$ by normalizing the frequency of each action over the total actions they performed during the roll-out. We measure specialization as the Jensen-Shannon divergence between $|\mathcal{N}|=2$ agents, $\text{JSD}(P_{1}||P_{2})$.
Task Parallelizability Prediction. Our model for predicting task parallelizability measures the relative efficiency gained from two generalist agents making soups in parallel versus specializing in subtasks (e.g., fetching ingredients versus serving soup). Consider the task of making one-onion soup. This requires $m=4$ subtasks: Onion $\to$ Pot, Cook in Pot, Bowl $\to$ Pot, and Pot $\to$ Serve. Each subtask $i$ is associated with (i) a number of workstations $\mathcal{B}_{r}(i)\in\mathcal{B}_{r}$, (ii) a path length $d(i)$, and (iii) a measure of spatial bottlenecks $\mathcal{B}_{s}(i)$ between those workstations, where $\mathcal{B}_{s}(i)\geq 0.8$ indicates constrained areas. If there are multiple paths, we consider $d_{\min}(i)$. Subtask durations are weighted by path length (Equation 5), with Cook assumed to take 10 time-steps. The overall speed-up expected from two generalist agents making soups in parallel is thus given by $S(N,\mathcal{B})$ in Equation 8. We hypothesize an inverse relationship between $S(N,\mathcal{B})$ and $\text{JSD}(P_{1}||P_{2})$. This would suggest agents specialize more when task and environmental constraints inhibit them from executing the same subtasks simultaneously.
4 Experiment 1: Exploring Specialization in Large-Scale Overcooked Simulations
Figure 3: Experiment 1 Results. (A) Histogram of specialization (JSD) in the best-performing seeds for 3,200 unique Overcooked configurations reveals a bimodal distribution, with teams clustering as either fully generalist or specialist. (B) Scatter plot of predicted task parallelizability ($S$) vs. observed JSD reveals a moderate negative correlation, indicating less specialization when a task has greater parallelizability. (C) Variables influencing parallelizability show strong, significant effects on JSD: specialization decreases with more pots but increases with greater spatial bottlenecks and task complexity. Error bars represent standard error.
In Experiment 1, we conduct a large-scale analysis across a variety of environments to test whether parallelizability impacts the emergence of specialization.
4.1 Experiment Design
We generated a test suite of 3,200 unique Overcooked environments for exploratory analysis. We utilized 14 spatial configurations representing 2D gridworld kitchens, each with a unique arrangement of counters and floor areas, including the original layouts from Overcooked-AI and a new set of layouts inspired by real-world kitchens (Pejic et al., 2019). A complete set of layouts is provided in Figure 6. To create diverse parallelizability constraints, we systematically varied the number of workstations (onions, tomatoes, pots, bowls, and serving stations). A total of 16 combinations included one or two onion piles, bowls, pots, serving stations, and zero or one tomatoes. Workstation positions were randomized three times per layout. Agents were trained to make five recipes: (i) 1-onion (ii) 2-onion (iii) 3-onion (iv) 1-onion and 1-tomato and (v) 2-onion and 1-tomato soup. These variations produced a total of 3,200 unique environments. For each of these environments, we trained two independent agents to play Overcooked collaboratively via self-play using PureJaxRL and PPO with five random seeds (Schulman et al., 2017; Lu et al., 2022). Full details regarding training procedures can be found in Appendix A.3. We then selected the highest-performing seed based on the joint reward that agents achieved during the roll-out. If multiple seeds reached the same maximum reward, we averaged their JSD. Trials were excluded if agents failed to take high-level actions or achieved a roll-out reward of 0.
4.2 Results
Distribution of JSD. Results from Experiment 1 are depicted in Figure 3. Our simulations revealed two distinct behavioral modes. Specialists ($JSD\geq 0.5$) dominated $74.9\%$ of runs, with $48.9\%$ fully specialized ($JSD>0.9$). Generalists ($JSD<0.5$) appeared in $25.1\%$ of runs, including $22.3\%$ fully generalist ($JSD<0.1$). The remaining $28.8\%$ exhibited mixed specialist-generalist behaviors ($0.1\leq JSD\leq 0.9$).
Effects of Parallelizability on Specialization. For each of the variables presented in our predictive model, there was a significant effect in the expected direction on observed JSD. Full statistical details can be found in Appendix A.4. We found a significant effect of the number of pots, but not the total number of workstations; spatial bottlenecks; and recipe complexity (the number of ingredients required to make a soup) on emergent specialization.
Model Evaluation. We analyzed the relationship between predicted task parallelizability and observed specialization (JSD) across 14 Overcooked layouts. Full statistical details are in Appendix A.5, with Tables 1 and 2 summarizing correlations, JSD, and rewards. The average correlation $-0.486$ ($p=8.49\times 10^{-28}$) revealed a moderate negative relationship between generalist efficiency and JSD, supporting our hypothesis that greater task parallelizability predicts lower specialization. A logistic regression classifying low vs. high specialization ($JSD\geq 0.5$) achieved $74.2\%$ accuracy, with $0.93$ precision ($F1=0.81$) for high specialization and $0.46$ precision ($F1=0.58$) for low specialization.
Effects of State Space on Specialization We also observed a moderate correlation between grid size and average JSD ($r=0.383,p=5.57\times 10^{-17}$), even when controlling for resource and spatial bottlenecks (Figure 7). This suggests that state space size and workstation distances may act as confounding variables, influencing specialization independently of task-specific constraints. Larger state spaces increase exploration demands. In response, agents may adopt specialization as a locally optimal strategy to reduce policy complexity by focusing on specific paths and workstation interactions.
5 Experiment 2: Validating Predictive Model
Figure 4: Experiment 2 Results. (A) Six layouts (four depicted) with varying bottlenecks were designed on a $5\times 5$ grid, each with one or two pots, two onions, one bowl, and one serving station across four locations, yielding 48 layouts. Each layout was paired with one-, two-, or three-onion soup recipes for a total of 144 trials. (B) Histogram of specialization (JSD) across the best-performing seeds reveals a distinctly bimodal distribution, with most clustering around low (generalist) or high (specialist) JSD. (C) Scatter plot of predicted task parallelizability ($S$) vs. observed JSD averaged. A strong negative correlation confirms that greater potential speed-up from generalists corresponds to lower specialization.
In Experiment 1, we confirmed that the variables which should affect generalist strategies from speeding up task completion have a significant impact on specialization. In Experiment 2, we validate our results using a smaller, controlled experiment. We also show that we can use our findings to design environments which will induce or inhibit the emergence of specialized roles.
5.1 Experiment Design
We utilized the same methods, game dynamics, and training schemes as in Experiment 1 (described in Section 4.1). To control state space size while varying spatial bottlenecks, we constructed six $5\times 5$ grid layouts: (1) open space with surrounding counters, (2) single vertical divider, (3) double vertical divider and (4) horizontal divider. These spatial layouts are illustrated in Figure 4. We varied the number of pots (1 or 2) rather than workstations, since pots had the largest effect on specialization in Experiment 1, while controlling for relative distances between workstations across trials. Each layout featured four workstation arrangements, totaling 48 configurations. For each, agents were trained to make one-, two-, or three-onion soup with 10 random seeds for a total of 1400 runs. As in Experiment 1, we again selected the highest-performing seed in terms of the joint reward that agents achieved during the policy roll-out after averaging over the randomized locations of workstations. Four trials were excluded from Experiment 2.
5.2 Results
Results from Experiment 2 are depicted in Figure 4. By fixing the size of the grid, the distinction between generalist and specialist strategies became more pronounced, with a clear bimodal distribution emerging. There were 81 specialist teams with $JSD\geq 0.5$ ($57.86\%$) and 59 generalist teams ($42.14\%$). Notably, the proportion of generalists increased substantially compared to Experiment 1 once we removed the confound of altering layout size, providing further evidence that specialization is partially driven by the size of the state space. Additionally, our predictive model achieved even greater accuracy for predicting specialization. The Pearson correlation between predicted task parallelizability and JSD was $-0.667$ ($p=8.94\times 10^{-6}$), confirming the hypothesized inverse relationship between generalist efficiency and specialization. Statistical details can be found in Appendix A.6. To validate the model’s predictive power, we fit a class-balanced logistic regression to classify high ($JSD\geq 0.5$) or low specialization ($JSD<0.5$). The model achieved a test set accuracy of $90.90\%$, demonstrating strong performance with $0.88$ precision ($F1=0.93$) for high specialization and $1.00$ precision ($F1=0.86$) for low.
These results suggest that spatial bottlenecks and resource availability are key factors driving the emergence of specialization. Our findings indicate that our predictive model can be used to systematically design environments that will either promote or inhibit specialization. Rather than tailoring training algorithms to induce policy diversity, it is possible to do so based on environmental design alone. If agents are given ample space and workstations to learn the task, they will develop emergent generalist behaviors. Alternatively, if the environment is constrained spatially or in terms of resources, agents will develop specialist roles.
6 Experiment 3: Investigating State Space Size and Specialization
In Experiment 3, we investigate the relationship between state space size and specialization in greater detail. Our findings show that as the state space expands, even if the size of the grid that the agents can explore is held constant, agents consistently converge on more specialized strategies, indicating that specialization is a less computationally costly alternative when learning is made more difficult. These findings highlight a challenge in MARL: as state spaces grow, agents struggle to converge on optimal generalist policies. Full details, including the experimental design, results and statistical analysis are provided in Appendix A.7 and A.8.
7 Related Work
The study of specialization in multi-agent systems spans multiple disciplines. Division of labor leading to functional specialization has been widely studied in biological and social domains, including cells (Bell & Mooers, 1997; Rueffler et al., 2012), insect colonies (Ratnieks & Anderson, 1999; Fjerdingstad & Crozier, 2006), and human societies (Kitcher, 1990; Kreider et al., 2022; Goldstone et al., 2024; Vélez et al., 2024). Inspired by these insights, research in MARL and robotics has sought to design algorithms that actively encourage specialization by promoting policy diversity (Murciano et al., 1997; Padgham & Winikoff, 2002; Zhu & Zhou, 2008; Charakorn et al., 2020; Wang et al., 2020a, b; Li et al., 2021; Bettini et al., 2023, 2024) or by organizing agents into hierarchies (Zhang et al., 2010; Ahilan & Dayan, 2019). Similarly, new work on language-agent collaboration often assumes that specialization is necessary, prompting agents to adopt distinct roles before collaborating (Wang et al., 2023; Suzgun & Kalai, 2024; Swanson et al., 2024; Liu et al., 2024; Juang et al., 2024). However, much of this work assumes that specialization is always beneficial, without systematically identifying when and why generalist strategies may outperform specialists. A useful contrast comes from distributed systems, which are explicitly engineered rather than emergent. In these systems, heterogeneity and parallelism are carefully designed to optimize performance (Moncrieff et al., 1996; Almasi & Gottlieb, 1994; Pancake, 1996) and minimize contention (Park & Reveliotis, 2001). This structured approach provides a useful reference point for understanding when emergent specialization is not just incidental, but optimal.
While biological and social systems provide compelling examples, it remains an open question when specialization emerges as the optimal architecture. Prior work suggests that under conditions of high environmental variability or individual failure, flexible generalist strategies may outperform rigidly specialized roles (Staps & Tarnita, 2022; Ben-Oren et al., 2023). In our work, we propose a novel theoretical framework for predicting the utility of specialization based on task efficiency, considering environmental bottlenecks, resource constraints, and task complexity. Prior work in RL has shown that spatial bottlenecks such as doorways can shape subgoals (Menache et al., 2002), while research in human-robot collaboration and MARL has demonstrated that environmental design influences emergent behavior (Fontaine et al., 2021; McKee et al., 2022). Building on these insights, we formalize a predictive model for when generalist versus specialist strategies should emerge.
8 Discussion
Specialization, or the differentiation of agents’ policies and roles, is often assumed to improve collaborative performance. However, principles from distributed systems offer an alternative hypothesis. Specifically, generalist agents may outperform specialist ones when concurrent execution of the entire task would improve speed and efficiency. In this work, we outlined a theoretical framework for predicting generalist versus specialist efficiency in cooperative multi-agent tasks inspired by a variant of Amdahl’s Law. We then validated this model across three Overcooked experiments, finding highly significant effects of parallelizability factors on emergent specialization and high predictive accuracy. Together, these findings suggest that specialization is constrained by the environment and parallelizability of a task, making it possible to design layouts which either encourage or inhibit role emergence regardless of whether the training algorithm explicitly incentivizes diversity. More broadly, this framework motivates the use of distributed systems in analyzing trade-offs that arise when dividing labor.
This work has several limitations. First, our model does not account for switching behavior (generalists who switch roles rather than cooking independently) or cases where specialist and generalist policies are equally optimal. This limitation arises from an implicit assumption from Amdahl’s Law that subtasks are executed sequentially without overlap. Second, our model provides a theoretical upper bound on the benefits of generalists, but does not account for potential coordination overhead when increasing the number of agents, which could make the empirical relationship non-monotonic. Third, our model assumes agents are equally capable, yet specialization can also arise when agents have heterogeneous skills that make them more efficient at certain tasks (Moncrieff et al., 1996; Ginting et al., 2022). Despite these limitations, our predictive model is highly extensible. It generalizes from 2 to $n$ agents, providing explicit predictions for larger teams. Finally, our work suggests that specialization is a locally optimal solution that agents use when exploration is costly, i.e. in larger state spaces. Clarifying this trade-off can provide insight into how specialization may improve sample efficiency, as well as how to decrease coordination overhead between agents.
Acknowledgments
This work was supported in part by the Department of Defense (DoD) through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program, the Office of Naval Research (ONR) MURI grant (N00014-22-1-2813), the EPSRC CDT in RAS (EP/L016834/1), and a grant from the Templeton World Charity Foundation.
Appendix A Appendix
A.1 Original Form of Amdahl’s Law
Let $f$ represent the proportion of a task that can be parallelized and accelerated by a factor of $s$, while $1-f$ denotes the portion that must be executed serially. The theoretical speedup $S$ achievable when distributing a task across multiple processors is given by:
$S=\frac{1}{(1-f)+\frac{f}{s}}$ (9)
A.2 Task Decomposition Graph
Figure 5: Example of a task graph for making one-onion soup in Overcooked. Task-relevant objects are represented as nodes in a directed acyclic graph (DAG), and edges represent subtasks.
A.3 Training Procedures
For each trial in our MARL experiments, we trained two independent agents to play Overcooked collaboratively through self-play. Our implementation builds upon the PureJaxRL framework for Proximal Policy Optimization (Schulman et al., 2017; Lu et al., 2022). Each agent was implemented using an soft Actor-Critic network (Haarnoja et al., 2018), in which both the actor and critic networks were composed of two fully-connected layers with 64 hidden units each. We trained each agent using Proximal Policy Optimization with an Adam optimizer and a learning rate of $2.5\times 10^{-3}$ for $1\times 10^{7}$ timesteps (Zhang, 2018). At each timestep, the agents observed the state of the current environment as a $\text{width}\times\text{height}\times 27$-dimensional grid. The first two dimensions corresponded to the spatial coordinates of the grid, while the 27-dimensional vector encoded the state of each grid cell, indicating whether it contained an agent, onion, tomato, bowl, or workstation. The policy produced a categorical distribution over discrete actions.
To accelerate learning in the sparse-reward environments, we applied reward shaping to augment raw rewards with an auxiliary-shaped signal. During reward shaping, each agent received rewards for performing specific actions aligned with the given recipe’s requirements: adding onions or tomatoes to the pot when an additional ingredient is needed, cooking the pot once it contains the correct quantity of ingredients, retrieving the soup, and finally serving it. The shaping reward $r_{s}$ was annealed over a predefined horizon ($2.5\times 10^{6}$) using a linear schedule.
A.4 Experiment 1: Significance Tests
Number of Pots and Workstations. First, we found a significant effect of the number of pots, but not the total number of workstations, on specialization. Agents tended to specialize more with one pot ($M=0.77,SD=0.30$) than two ($M=0.57,SD=0.41$) but not when presented with more of other workstations (e.g., bowls, serving stations). A Welch’s t-test between layouts with one pot versus two revealed a significant difference in JSD, $t(2842.72)=14.98,p=7.23\times 10^{-49}$. Conversely, a one-way ANOVA showed a non-significant difference in JSD across the full range of possible workstation combinations ($F(5,3094)=0.82,p=0.54$).
Spatial Bottlenecks. Next, we found a significant effect of spatial bottlenecks on specialization. Agents specialized less in layouts with fewer bottlenecks ($M=0.61,SD=0.38$) than in those with medium ($M=0.69,SD=0.37$) or high bottlenecks ($M=0.86,SD=0.27$). Using edge betweenness centrality, we quantified the average extent of spatial bottlenecks across the entire layout. Betweenness values were divided into three equal-width bins: Low $[2.110,3.194)$, Medium $(3.195,4.277]$, and High $(4.277,5.360]$. Smaller centrality values represented a fewer number of nodes that were critical to many paths in the layout graph, corresponding to a smaller number of spatial bottlenecks. Larger centrality values represented more critical nodes, corresponding to greater spatial bottlenecks. A one-way ANOVA showed a significant difference in JSD across bins ($F(2,2097)=34.10,p=2.26\times 10^{-15}$). Pairwise comparisons using Tukey’s HSD showed that JSD was significantly higher with medium centrality than low ($p=3.18\times 10^{-7}$), high centrality than medium ($p=1.07\times 10^{-12}$), and high centrality than low ($p=3.82\times 10^{-20}$).
Recipe Complexity. Finally, we found a significant effect of recipe complexity (the number of ingredients required to make a soup) on JSD. Agents tended to specialize less when recipes had one ingredient ($M=0.54,SD=0.47$) than with two ($M=0.73,SD=0.35$) or three ($M=0.67,SD=0.32$). A one-way ANOVA showed a significant difference in JSD across recipes ($F(2,3097)=57.22,p=3.95\times 10^{-25}$). Pairwise comparisons using Tukey’s HSD revealed that JSD was significantly higher for recipes with two ingredients compared to one ($p=4.46\times 10^{-19}$) and recipes with three ingredients compared to one ($p=4.37\times 10^{-10}$), but significantly lower for recipes with three ingredients compared to two ($p=7.77\times 10^{-6}$).
A.5 Experiment 1: Evaluation
For each of the 14 layouts, there were a total of 16 workstation combinations with three randomized positions each. We computed the predicted parallelizability, $S$, for each of these combinations and averaged over the three randomized locations to compute a measure of task parallelizability for each combination of layout, workstations, and recipe.
We also fit a logistic regression model where the binary outcome variable was low versus high specialization (JSD $\geq 0.5$) and the predictor variable was task parallelizability. To address the class imbalance in the JSD distribution, we used class-weight balancing to automatically adjust class weights inversely proportional to their frequencies in the training data.
Figure 6: Overcooked Spatial Layouts in Experiment 1.
Figure 7: Effect of grid size on observed specialization (JSD) in Experiment 1. As the size of the grid increased, agends tended to specialize more (even when presented with the same number of workstations and bottlenecks). W
Spatial Layout Mean JSD Mean Reward
H_shape 0.666725 14.165217
L_shaped 0.599760 14.050847
U_shaped 0.588023 15.222689
big_room 0.732749 13.207627
bottleneck_room 0.686606 14.212121
coord_ring 0.587970 14.845188
counter_circuit 0.829034 12.990385
cramped_room 0.597489 15.879167
gallery 0.603094 14.564854
semi_forced_coord 0.744749 13.440191
signaling_room 0.860218 12.813953
single_wall 0.628655 12.808696
too_many_cooks 0.631687 14.099567
two_islands 0.758321 13.133005
Table 1: Average specialization (JSD) observed in different spatial layouts in Experiment 1.
Workstations Mean JSD Std JSD
1 onions 0.69 0.38
2 onions 0.66 0.40
1 pots 0.76 0.33
2 pots 0.58 0.42
1 serving 0.66 0.39
2 serving 0.69 0.38
0 tomatoes 0.60 0.41
1 tomatoes 0.79 0.31
1 bowls 0.68 0.38
2 bowls 0.67 0.39
Table 2: Mean and standard deviation of observed specialization (JSD) for varying numbers of workstations in Experiment 1.
Recipes Mean JSD Std JSD
1 onion 0.56 0.48
2 onions 0.65 0.39
3 onions 0.59 0.36
1 onion 1 tomato 0.82 0.33
2 onions 1 tomato 0.77 0.29
Table 3: Mean and standard deviation observed specialization (JSD) for different recipes in Experiment 1.
Table 4: Correlations Across Spatial Layouts
Base Layout Correlation p-value
H_shape -0.493 0.004
L_shaped -0.699 $<0.001$
U_shaped -0.679 $<0.002$
big_room -0.568 0.001
bottleneck_room -0.544 0.001
coord_ring -0.307 0.087
counter_circuit -0.361 0.042
cramped_room -0.619 $<0.001$
gallery -0.524 0.002
semi_forced_coord -0.555 0.001
signaling_room -0.479 0.009
single_wall -0.553 0.001
too_many_cooks -0.272 0.132
two_islands -0.244 0.178
Average Correlation -0.494
Table 5: Correlations between predicted task parallelizability ($S$) and observed specialization (JSD) across varying spatial layouts in Experiment 1. The significance of this relationship varied across layouts, especially those with different grid sizes, prompting the controlled design of Experiments 2 and 3.
Figure 8: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific spatial layout.
Figure 9: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution for a specific number of workstations.
Figure 10: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific recipe.
Figure 11: Specialization (JSD) distributions across all random seeds in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific layout. The distributions of strategies appear mostly consistent across different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
Figure 12: Specialization (JSD) distributions across all random seeds in Experiment 1. Each subfigure represents the JSD distribution grouped by the number of workstations. The distributions appear consistent across the different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
Figure 13: Specialization (JSD) distributions across all random seeds in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific recipe. The distributions appear consistent across the different random seeds. Five random seeds were used: 1, 2, 3, 5, and 7.
A.6 Experiment 2: Results
Number of Pots Mean JSD Mean Reward
1 0.676053 14.861111
2 0.280481 18.867647
Table 6: Average specialization (JSD) with varying numbers of pots in Experiment 2.
Spatial Layout Mean JSD Mean Reward
Open 0.3320525 18.7083330
Horizontal Divider 0.332101 18.0416668
Vertical Divider 0.318273 18.0000000
Vertical Space 0.404964 18.0833335
Long Divider 0.752011 12.2083334
Two Dividers 0.2988275 17.6666666
Table 7: Average specialization (JSD) in various spatial layouts in Experiment 2.
A.7 Experiment 3: Design
Figure 14: Three layouts from Experiment 3 are shown: (a) the smallest layout, (b) the largest layout with increased exploration, and (c) the largest layout with consistent exploration. Across these configurations, the optimal policy remains constant. The plots depict how the JSD and reward vary with increasing layout size under two different conditions. In the case of the three-onion soup, the reward decreases as the layout size grows. The smallest layout (layout 1) corresponds to the minimum JSD and the maximum reward.
In Experiment 3, we examined the influence of various divider configurations on agents specialization. Four skeleton layouts of identical size were tested: (1) a single vertical divider, (2) a double vertical divider, (3) a horizontal divider, and (4) no divider. The focal point was to investigate how these structural variations affect the convergence of agent strategies. To evaluate the impact of state space size on optimal policy convergence and in cases of suboptimal convergence, whether agents developed more specialist or generalist strategies - we conducted a controlled experiment comprised of 280 runs. The experiment systematically varied seven layout sizes, utilizing 10 random seeds per layout, two recipe configurations (1-onion soup and 3-onion soup), and two experimental conditions (constant vs. non-constant exploration). Critically, the optimal policy remained consistent across all runs; only empty space that the agents could not access was added to increase the size of the state space. Figure 14 illustrates three of these layouts, demonstrating how variations in layout size were implemented, while keeping the optimal policy consistent
A.8 Experiment 3: Results
For the three-onion soup task, the majority of run converged to a JSD value fo approximately 0.6 - achieved in $93.8\%$ of runs with a single divider, $87.5\%$ of runs with a double divider, $87.5\%$ in the no-divider room layout, and $93.8\%$ in the counter circuit layout. Notably, a JSD value of 0 corresponds to the optimal strategy - however, agents rarely reached this level of convergence. For the one-onion soup tasks, full specialization was observed in most runs across the four layouts; $100\%$ with a single divider, $81.3\%$ with a double divider, and $81.3\%$ with no-divider. In stark contrast, only $46.3\%$ of runs in the counter circuit layout converged to a specialist strategy ($JSD>0.5$), while the remaining $53.4\%$ of runs resulted in generalist behavior.
The differences in emergent strategies between layouts is likely due to the distinct effects of the dividers on agent behavior. In layouts with vertical dividers, agents are more inclined to specialize, as the dividers channel them toward the same region where they cannot work in parallel. Conversely, the counter circuit layout encourages generalist strategies by focusing them to either occupy separate regions or follow each other, reducing the likelihood of specialization.
The effect of layouts size varied between the two recipes. For the simpler 1-onion soup recipe, there was no significant relationship between layout size and either reward JSD. Under non-constant exploration, JSD increased significantly with layout size ($r=0.7924,p=0.0336$) indicating greater policy specialisation in large layouts, but reward showed no reliable trend ($r=-0.2512,p=0.5869$). Pairwise T-tests confirmed no significant difference between the smallest and largest environments for JSD ($p=0.755$ or reward $p=0.210$. Under constant exploration, the 1-onion soup was unaffected by layout size, with no substantial correlation for JSD ($r=-0.5850,p=0.1677$) or reward ($r=0.0609,p=0.8969$).
In contrast, the 3-onion soup recipe was sensitive to increasing layout size. Under non-constant exploration, as layout size increased, JSD rose significantly ($r=0.7549,p=0.0498$) while reward decreased ($r=-0.8556,p=0.0140$), indicating that agents struggle to perform well in large environments and develop more specialised policies. Pairwise T-tests further outlined significant differences between the smallest and largest layouts for JSD ($p=0.008$ and reward $p<10^{-4}$. Even under constant exploration, the 3-onion soup recipe showed a decline in reward as layout size increased ($r=-0.4308,p=0.0020$), although JSD was not significantly affected.
These findings demonstrate that increasing both layout size and the number of subtasks, thus expanding the exploration space, drives agents toward suboptimal and overly specialized policies. Even when exploration remains constant, the addition of extraneous state dimensions in larger layouts prevents the RL policy’s ability to converge effectively to the optimal solution. This underscores a key challenge for future work: MARL policies struggle to consistently identify optimal generalist solutions in larger layouts, even when task structure, workstations, and exploration dynamics remain unchanged.
Layout Experiment Recipe Mean JSD Std JSD Count Mean Reward Std Reward Reward Count
1 A 1-Onion 0.750219 0.397342 10 24.8 6.014797 10
1 A 3-Onion 0.307466 0.477938 10 13.7 3.301515 10
1 B 1-Onion 0.750219 0.397342 10 24.8 6.014797 10
1 B 3-Onion 0.307466 0.477938 10 13.7 3.301515 10
2 A 1-Onion 0.570917 0.481485 10 24.0 4.136558 10
2 A 3-Onion 0.712140 0.355738 10 9.8 3.047768 10
2 B 1-Onion 0.751195 0.431745 9 27.8 7.052186 10
2 B 3-Onion 0.991283 0.013504 6 8.4 0.516398 10
3 A 1-Onion 0.683808 0.453647 10 27.2 5.329165 10
3 A 3-Onion 0.428280 0.458991 9 9.5 4.478343 10
3 B 1-Onion 0.579108 0.497131 10 28.9 6.136412 10
3 B 3-Onion 0.887113 0.197573 10 8.7 0.483046 10
4 A 1-Onion 0.956536 0.130393 9 23.2 8.202980 10
4 A 3-Onion 0.686612 0.392708 6 8.0 3.829708 10
4 B 1-Onion 0.462766 0.497691 10 31.3 4.922736 10
4 B 3-Onion 0.657041 0.433589 5 8.2 3.794733 10
5 A 1-Onion 0.806959 0.310521 10 24.2 3.119829 10
5 A 3-Onion 0.851840 0.235864 6 6.7 5.355164 10
5 B 1-Onion 0.694087 0.476603 10 28.9 4.931757 10
5 B 3-Onion 0.478804 0.391161 6 7.9 3.348300 10
6 A 1-Onion 0.975668 0.072995 9 24.2 3.293090 10
6 A 3-Onion 0.899543 0.210584 7 8.3 0.483046 10
6 B 1-Onion 0.504944 0.451051 10 26.4 11.393956 10
6 B 3-Onion 0.520999 0.319708 7 9.4 2.988868 10
7 A 1-Onion 1.000000 0.000000 8 24.3 3.335000 10
7 A 3-Onion 0.750629 0.354937 8 6.7 3.560587 10
7 B 1-Onion 0.583097 0.488816 8 26.3 7.424434 10
7 B 3-Onion 0.713627 0.403798 5 6.5 3.439961 10
Table 8: Detailed results from Experiment 3. Each row represents a unique combination of spatial layout, experiment, and recipe, showing the mean, standard deviation (std), and count (number of runs) for both specialization (JSD) and reward metrics. Experiment A corresponds to expanding exploration, while Experiment B denotes constant exploration. The JSD count is less than 10 for certain permutations because the policy failed to converge, resulting in the agents taking no meaningful actions.
Recipe Exploration Metric Corr. p-value
1-Onion Expanding JSD 0.7924 0.0336
Reward -0.2512 0.5869
Constant JSD -0.5850 0.1677
Reward 0.0609 0.8969
3-Onion Expanding JSD 0.7549 0.0498
Reward -0.8556 0.0140
Constant JSD -0.0423 0.9283
Reward -0.6947 0.0832
Table 9: Correlations from Experiment 3. Correlations between specialization (JSD) and reward for one-onion and three-onion Soup recipes under expanding and constant exploration conditions. The table reports correlations and p-values ($p$) for JSD and reward in relation to layout size.
Condition Metric t-value p-value
Combined Recipes
layout_1 vs layout_7 JSD -2.229 0.029
layout_0 vs layout_7 Reward 0.506 0.614
1-Onion Soup Recipe
layout_1 vs layout_7 JSD -0.315 0.755
layout_1 vs layout_7 Reward -1.278 0.210
3-Onion Soup Recipe
layout_1 vs layout_7 JSD -2.820 0.008
layout_1 vs layout_7 Reward 5.812 0.000
Table 10: Results from Experiment 3. Pairwise T-tests comparing the smallest (Layout 1) layout with the largest layout (Layout 67).

### v2

@@
 Abstract
-Multi-agent systems often rely on specialized agents with distinct roles rather than general-purpose agents that perform the entire task independently.
-However, the conditions that govern the optimal degree of specialization remain poorly understood.
-In this work, we propose that specialist teams outperform generalist ones when environmental constraints limit task parallelizability—the potential to execute task components concurrently.
-Drawing inspiration from distributed systems, we introduce a heuristic to predict the relative efficiency of generalist versus specialist teams by estimating the speed-up achieved when two agents perform a task in parallel rather than focus on complementary subtasks.
-We validate this heuristic through three multi-agent reinforcement learning (MARL) experiments in Overcooked-AI, demonstrating that key factors limiting task parallelizability influence specialization.
-We also observe that as the state space expands, agents tend to converge on specialist strategies, even when generalist ones are theoretically more efficient, highlighting potential biases in MARL training algorithms.
-Our findings provide a principled framework for interpreting specialization given the task and environment, and introduce a novel benchmark for evaluating whether MARL finds optimal strategies.
-multi-agent reinforcement learning, collaboration, efficiency, specialization, policy diversity
+When should we encourage specialization in multi-agent systems versus train generalists that perform the entire task independently?
+We propose that specialization largely depends on task parallelizability: the potential for multiple agents to execute task components concurrently.
+Drawing inspiration from Amdahl’s Law in distributed systems, we present a closed-form bound that predicts when specialization improves performance, depending only on task concurrency and team size.
+We validate our model on two standard MARL benchmarks that represent opposite regimes— StarCraft Multi-Agent Challenge (SMAC, unlimited concurrency) and Multi-Particle Environment (MPE, unit-capacity bottlenecks)—and observe close alignment between the bound at each extreme and an empirical measure of specialization.
+Three follow-up experiments in Overcooked-AI demonstrate that the model works in environments with more complex spatial and resource bottlenecks that allow for a range of strategies.
+Beyond prediction, the bound also serves as a diagnostic tool, highlighting biases in MARL training algorithms that cause sub-optimal convergence to specialist strategies with larger state spaces.
 1 Introduction
 Collaboration enables multi-agent systems to achieve goals that exceed the capabilities of a single agent.
-Advances in artificial intelligence (AI) have enhanced the capabilities of multi-agent systems to complete complex, coordinated tasks, including automating warehouses (Canese et al., 2021), sustainable resource management (Perolat et al., 2017), and simulated social interactions (Park et al., 2023).
-However, efficient task allocation remains a key challenge.
-Existing work across multiple fields, including multi-agent reinforcement learning (MARL), cognitive science, and robotics, often assumes that specialization is universally desirable (Padgham & Winikoff, 2002; Zhu & Zhou, 2008; Charakorn et al., 2020; Wang et al., 2020a, b; Li et al., 2021; Bettini et al., 2023; Juang et al., 2024; Bettini et al., 2024; Swanson et al., 2024).
+Advances in artificial intelligence (AI) have enhanced the capabilities of multi-agent systems to complete complex, coordinated tasks, including automating warehouses [1], sustainable resource management [2], and simulated social interactions [3].
+The prevailing assumption in multi-agent reinforcement learning (MARL), cognitive science, and robotics is that specialization is universally desirable [4, 5, 6, 7, 8, 9, 10, 11, 12, 13].
 For example, fine dining restaurants rely on specialists—the saucier prepares sauces, the rôtisseur roasts, and so on—who each perform distinct yet complementary subtasks.
-Intuitively, specialization narrows action spaces (Wang et al., 2020b), reduces ambiguity about what agents should do next (Goldstone et al., 2024), and eases individual cognitive loads (Griffiths, 2020).
-This frames task allocation as an optimization problem, focusing on identifying roles and assigning agents accordingly (Wang et al., 2020a).
-However, human teams demonstrate that specialization is not always optimal.
-Fine dining chefs coordinate to produce complex meals, but at the cost of high training overhead, rigidity, and failure risk if a key specialist is unavailable (Ben-Oren et al., 2023; Vélez et al., 2024).
-In contrast, generalist fry cooks perform the same subtasks in parallel, such as grilling and assembling, maximizing adaptability and throughput.
-Examples of specialist and generalist strategies are depicted in Figure 1.
-Furthermore, although AI agents also face coordination challenges (Shoham & Leyton-Brown, 2008; Wu et al., 2021), it remains unclear whether the constraints that make specialization advantageous for people also apply.
-With agents becoming increasingly capable we must revisit this assumption and ask: under what conditions, if any, is specialization truly optimal?
+Intuitively, specialization narrows action spaces, reduces ambiguity about what agents should do next [14], and eases individual cognitive loads [15].
+Recent algorithms such as ROMA and RODE rely on this assumption to treat task allocation as an optimization problem, focusing on learning roles and constraining agents to specific subtasks [8, 7].
+However, in both human and MARL teams, specialization introduces trade-offs.
+Fine dining chefs produce complex meals via rigid role divisions, but the system is brittle: training costs are high, and service halts if one specialist fails or takes longer than the others [16, 17].
+In contrast, generalist fry cooks, who all grill, assemble, and plate, maximize throughput by performing subtasks in parallel.
+While MARL agents do not fatigue or quit like humans, they do face coordination challenges, policy divergence, and retraining costs when roles are overly fixed [18, 19].
+This raises a key question: what task and environmental constraints make specialization truly optimal, and when are generalists better?
 Figure 1: Specialists vs.
 generalists.
-(A) Specialist teams tend to focus on non-overlapping subtasks.
-In Overcooked, an example of this is Agent 1 putting onions and tomatoes in the pot, while Agent 2 fetches a bowl and serves the soup.
-In highly specialized teams, agents’ high-level action distributions are completely distinct, resulting in a Jensen-Shannon divergence ($\text{JSD}(P_{1}||P_{2})$) of 1.0.
-Example specialist video.
-(B) In generalist teams, both agents perform all subtasks.
-In Overcooked, this may look like Agents 1 and 2 independently making their own soups in parallel.
-In generalist teams, agents’ action distributions are the same, resulting in $\text{JSD}(P_{1}||P_{2})=0.0$.
-Example generalist video.
-We propose that a task’s parallelizability—the potential for its independent subtasks to be executed simultaneously—limits the benefits of generalist teams.
-Drawing from the distributed computing literature (Almasi & Gottlieb, 1994; McCool et al., 2012; Vélez et al., 2023), we use Amdahl’s Law to determine when multi-agent systems should favor generalists or specialists.
-Amdahl’s Law predicts that there are diminishing returns to increasing the number of processors performing a task in parallel, based on the portion of the task that must be performed serially (Amdahl, 1967; Hill & Marty, 2008; Cassidy & Andreou, 2011).
-Applied to multi-agent systems, this places an upper bound on performance gains achievable from generalist agents working in parallel.
-When spatial and resource constraints limit parallel execution, agents should improve efficiency by specializing in complementary subtasks.
-Using a simple, interpretable model, we provide a novel framework for evaluating the trade-offs between specialist and generalist strategies.
-To test this prediction, we develop a theoretical framework for comparing generalist and specialist efficiency and apply it to three multi-agent experiments in Overcooked-AI.
-Experiment 1 shows that key model variables influence specialization, with a moderate negative correlation between predicted task parallelizability and specialization, though larger state spaces complicate this relationship.
-Experiment 2 controls for these confounds, demonstrating our model’s high accuracy in predicting specialization.
-Experiment 3 finds that larger state spaces promote specialization by increasing coordination challenges.
-Our findings suggest that specialization is driven by task parallelizability and environment, enabling the design of layouts that encourage or inhibit policy diversity.
-More broadly, our framework highlights the value of using insights from distributed computing to improve the interpretability of multi-agent systems.
-2 Theoretical Framework
-We introduce a general framework for evaluating task parallelizability and predicting emergent specialization in multi-agent systems.
+(A) Specialist teams focus on non-overlapping subtasks, resulting in high Jensen-Shannon divergence (SI) between agents’ action distributions.
+Example video.
+(B) In generalist teams, both agents perform the task independently, resulting in low SI.
+Video.
+We propose that the degree to which a task can be executed in parallel (or its parallelizability) directly governs the effectiveness of generalist teams.
+When spatial or resource bottlenecks force agents to wait, the benefits of independently performing a task in parallel vanish and specialist policies become more efficient.
+Drawing from distributed computing [20, 21], we adapt Amdahl’s Law to treat agents as processors and subtasks as units of work that may or may not be parallelizable [22, 23, 24].
+This yields a simple analytic bound on the speed-up achievable by generalist teams as a function of task structure.
+When the parallelizability of the overall task falls below team size, specialization is advantageous.
+We show that a task’s parallelizability provides a simple prediction for when specialist policies will emerge in multi-agent systems given a team size and environment.
+In Experiment 1, we apply the model to two edge-case MARL benchmarks: SMAC, where generalists dominate due to unlimited concurrency, and MPE, where unit-capacity bottlenecks force agents to specialize.
+In Experiments 2 and 3, we extend to more complex spatial and resource dynamics in Overcooked-AI, demonstrating a more fine-grained relationship between task parallelizability and specialization.
+Experiment 4 shows that larger state spaces promote deviations from model predictions, highlighting biases in existing MARL training regimes.
+Together, these findings suggest that specialization is not obligatory for successful teams, and is affected by bottlenecks and coordination structure.
+Our framework enables the design of environments that encourage or inhibit policy diversity and illustrates how tools from distributed systems can illuminate multi-agent settings.
+2 Task Parallelizability Framework
+In many multi-agent settings, the marginal benefit of adding agents is limited not only by training but also structural constraints on the task and environment.
+Quantifying task parallelizability answers two key questions: (1) When should we encourage behavioral diversity and specialization during training, and (2) When parallelism is limited, which dimension— team size $N$, subtask allocation $f_{i}$, spatial capacity $C^{s}$, or resource capacity $C^{r}$— is responsible for preventing speed-up?
+The key result (§2.4) shows that when overall parallelizability $S(N,B)$ drops below team size, specialization becomes strictly more efficient.
+Sections 3–5 demonstrate how to measure concurrency constraints for various MARL applications and validate the prediction empirically.
 2.1 $N$-Player Markov Game
-Following Carroll et al.
-(2019) and Ruhdorfer et al.
-(2024), we formalize a cooperative decentralized Markov decision process (MDP), defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{N},A,\mathcal{T},\mathcal{R},\gamma\rangle$.
+Following Albrecht et al.
+[25], we formalize a cooperative decentralized Markov decision process (Dec-MDP), defined as $\mathcal{M}=\langle\mathcal{S},\mathcal{N},A,\mathcal{T},\mathcal{R},\gamma\rangle$.
 $\mathcal{S}$ is a finite set of states, and $\mathcal{N}$ is a set of $N$ agents.
@@
 Finally, $\mathcal{R}:\mathcal{S}\times A\to\mathbb{R}$ is the shared reward function.
-The objective of each agent is to maximize its expected discounted return $\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\big{]}$, where $\gamma\in[0,1)$ is a discount factor and $r(s_{t},a_{t})$ is the shared reward received by all agents at time $t$.
-While agents share the same reward, each optimizes its own cumulative return independently.
+Each agent’s objective is to maximize its expected discounted return $\mathbb{E}\big[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\big]$, where $\gamma\in[0,1)$ is a discount factor and $r(s_{t},a_{t})$ is the shared reward received by all agents at time $t$.
+Agents share the same reward but optimize their returns independently.
 2.2 Jensen-Shannon Divergence and Specialization
@@
 Let $\pi^{i}:\mathcal{S}\to\Delta(\mathcal{A}^{i})$ be the policy of agent $i$ that maps each state $s\in\mathcal{S}$ to a distribution over actions $a\in\mathcal{A}^{i}$.
-In generalist teams, agents do not specialize in specific subtasks, and thus select from the entire set of actions $A^{i}$.
-Formally, $\forall a\in\mathcal{A}^{i},\pi_{generalist}^{i}(a|s)>0\text{ for some }s\in\mathcal{S}$.
-In specialist teams, agents are restricted to specific subtasks and resources.
-In this case, $\mathcal{A}_{specialist}^{i}\subset\mathcal{A}^{i}$, and their specialist policies satisfy $\forall a\notin\mathcal{A}_{specialist}^{i},\pi_{specialist}^{i}(a|s)=0$.
+Each agent follows a state-visitation distribution that describes how frequently they occupy different states and actions under $\pi^{i}$, which can be used to define their effective action distributions $P_{i}(a)$ [26].
+For a complete formulation, refer to Appendix A.
+We approximate specialization using the Jensen-Shannon divergence (JSD) over agents’ action distributions $P_{i}$ [27, 28].
+We use JS rather than Kullback-Leibler divergence because it is (a) well-defined when agents’ action distributions have disjoint support and (b) symmetric ($\text{SI}(P_{1}||P_{2})=\text{SI}(P_{2}||P_{1})$), since the degree of specialization should be the same between teams regardless of the order of agents.
+For $N$ agents with action distributions $P_{1},P_{2},...,P_{N}$, JSD is given by:
+$\operatorname{JSD}(P_{1:N})\;=\;\frac{1}{N}\sum_{i=1}^{N}D_{\mathrm{KL}}\!\bigl(P_{i}\,\|\,M\bigr),\qquad M=\frac{1}{N}\sum_{i=1}^{N}P_{i},$ (1)
+where $M$ is the mixture distribution and $D_{\mathrm{KL}}(P_{i}||M)$ is the Kullback-Leibler (KL) divergence between each agent’s action distribution $P_{i}$ and $M$.
+Because the maximal value grows as $\max\operatorname{JSD}(P_{1:N})=\log_{2}N$ bits, we report the Specialization Index:
+$\mathrm{SI}(P_{1:N})\;=\;\frac{\operatorname{JSD}(P_{1:N})}{\log_{2}N}\;\in[0,1]$ (2)
+throughout the paper.
+$\mathrm{SI}=0$ indicates identical distributions; $\mathrm{SI}=1$ indicates perfectly non‑overlapping distributions.
+In fully generalist teams, agents select from the entire set of actions $A$ to complete the task in parallel.
+In this case, all $N$ agents have performed the same proportions of actions, leading to $\mathrm{SI}(P_{1},P_{2},...,P_{N})=0.0$ as an indicator of no specialization.
+Alternatively, in fully specialist teams, the action distributions of each agent are completely disjoint (i.e., each agent assigns probability exclusively to a unique set of actions), meaning that $\mathrm{SI}(P_{1},P_{2},...,P_{N})=1.0$ indicates complete specialization.
+Intermediate values of SI indicate partial specialization, in which the agents overlap in some subtasks or occupancies but not others.
 Examples of action distributions and specialist versus generalist strategies can be found in Figure 1.
-Let the set of high-level actions be denoted as $\mathcal{A}=\{a_{1},a_{2},...,a_{k}\}$ (Amato et al., 2019).
-Each agent follows a state-visitation distribution that describes how frequently they occupy different states under their policy (Schulman et al., 2015) given by:
+2.3 Multi-Agent Bottlenecks
+In distributed computing systems, Amdahl’s Law provides an upper bound on the speedup achievable through parallelism based on the proportion of a task that must be performed sequentially by a single processor [22, 23, 24].
+Certain computations inherently depend on centralized coordination or the results of previous steps [20, 21, 29].
+A classic example is memory bandwidth in GPUs, where increasing the number of threads does not lead to proportional speed-up due to limited shared access [30].
+By extending Amdahl’s Law to multi-agent systems, we can formulate a principled prediction for when tasks can be performed concurrently by agents or must be performed sequentially by one.
+We begin by identifying bottlenecks that prevent parallelism in multi-agent tasks.
+Various task constraints have been identified across numerous multi-agent domains, such as warehouse operations (insufficient space, storage locations, and pallets [31]), autonomous vehicles (lane reductions, traffic congestion [32]), and smart factories (equipment, production routes) [33].
+Inspired by these findings, we account for two key types of bottlenecks: (1) spatial bottlenecks limit the number of agents ($C^{s}$) that can navigate due to obstacles, narrow passages, or congestion, and (2) resource bottlenecks limit accesses ($C^{r}$) that can be made at a time to critical resources such as workstations or tools [34].
+Task decomposition.
+Below, we offer an operationalization of bottlenecks that can be applied to a wide variety of tasks.
+Following Solway et al.
+[35] and Amato et al.
+[36], we treat a cooperative task as a directed acyclic graph (DAG) $T=(V,E)$.
+Each node $i\in V$ is a subtask that can be executed by one or more agents, and an edge $(i\!\to\!j)\in E$ indicates that $j$ cannot start until $i$ completes.
+Let $m=|V|$ and denote by $f_{i}$ the expected fraction of total execution time spent in node $i$.
+We require only $\sum_{i=1}^{m}f_{i}=1$; concrete estimators for $f_{i}$ are introduced per domain in §3.
+Concurrency capacity.
+For every subtask $i$, let the concurrency capacity $C_{i}=\min\bigl(C^{s}_{i},\,C^{r}_{i}\bigr)$ denote the maximum number of agents that can execute subtask $i$ in parallel before contention forces at least one agent to wait.
+$C^{s}_{i}$ captures purely spatial limits (e.g.
+a one-cell doorway), while $C^{r}_{i}$ captures resource limits (e.g.
+two welding stations).
+We define subtasks independently of spatial location; two agents chopping vegetables on opposite ends of the room are treated as executing the same subtask.
+Although the exact estimator is environment-specific (see §3 for examples), most simulators support enumerating access paths or available resources and computing path widths.
+Alternatively, it is possible to count simultaneous agent accesses to each subtask during a short probe rollout.
+The minimum of these counts yields $C_{i}$.
+While we focus on spatial layout and resource access as primary limits of concurrency, other environments might impose additional bottlenecks, such as the latency required to exchange messages in communication-based environments.
+These bottlenecks, though not explored in this paper, can also be integrated into this concurrency score.
+2.4 Task Parallelizability Prediction
+We now adapt Amdahl’s Law to multi-agent tasks.
+For the original form, see Appendix B.
+Mieczkowski et al.
+[34] extended this idea to predict idleness during human collaborations given the current group size, workload, and resource constraints.
+We hypothesize that a similar formulation can be used to predict parallelizability in MARL.
+Let $f_{i}$ represent the fraction of the total task time required for the $i^{th}$ of $m$ subtasks such that $\sum_{i=1}^{m}f_{i}=1$.
+For each subtask $i$, the overall bottleneck capacity is defined as the minimum bottleneck that prevents agents executing it concurrently, or $C_{i}=\min(C^{s}_{i},C^{r}_{i})$.
+The subtask speed-up factor $s_{i}(N,C_{i})$ quantifies the efficiency gain from adding $N$ agents to a subtask, as opposed to a single agent performing it, given that $C_{i}$ bottlenecks restrict parallel execution.
+Formally, $s_{i}(N,C_{i})=\text{min}(N,C_{i})$.
+Thus, if we assume that there is no cost to switching between subtasks or coordinating with larger teams, the speed-up achievable when $N$ agents perform the entire task in parallel, which we call task parallelizability, is expressed as:
+$S(N,C)=\frac{1}{\sum_{i=1}^{m}\frac{f_{i}}{s_{i}(N,C_{i})}}$ (3)
+Properties
+There are several useful properties of this speed-up bound $S(N,C)$ provided in Appendix C.
+These propositions formalize the intuition that removing bottlenecks promotes generalists, never specialists, since only generalists benefit from concurrent access to subtasks.
+3 Applications to MARL Environments
+The bound in §2 is abstract; it relies only on the subtask breakdown $f_{i}$ and the concurrency capacity $C_{i}$.
+Here we show how to measure these values in three representative classes of multi-agent benchmarks: grid-worlds (Overcooked-AI),continuous environments (MPE), and high-throughput RTS scenarios (SMAC).
+In each domain, we define subtasks, estimate $f_{i}$, compute spatial and resource capacities, and plug them into Equation 3.
+StarCraft Multi-Agent Challenge (SMAC)
+In SMAC [37], agents cooperate to defeat enemy units.
+Each agent observes a local feature vector and selects from discrete actions such as left or attack enemy 1.
+Let $m$ (subtasks) be the number of enemy units that agents must defeat to gain reward.
+We assume all subtasks require roughly equal effort and set $f_{i}=\frac{1}{m}\text{ }\forall i\in m$.
+Since all agents can move and shoot freely, and combat maps are open with no obstacles, there are no spatial or resource bottlenecks.
+We define $C_{i}=\infty\text{ }\forall i\in m$.
+Substituting into Equation 3, we obtain $S(N,B)=N$, or maximum parallelizability.
+According to Proposition 2, we would predict fully generalist strategies with near-zero specialization (SI =0=0).
+Multi-Particle Environment (MPE)
+In MPE’s Spread task [38], agents must each cover a landmark.
+Let $m$ be the number of landmarks, and assume $f_{i}=1/m$.
+Since only one agent can occupy each landmark at a time, $C_{i}=1$ for all $i$, yielding $S(N,B)=1$.
+This low parallelizability should lead to full specialization (SI =1=1).
+Overcooked-AI
+In Overcooked-AI [39, 19], agents must cook recipes by navigating a grid and coordinating at shared workstations.
+We represent each recipe as a directed acyclic graph (DAG) of subtasks.
+For example, preparing onion soup involves the sequence: Onion → Pot → Cook → Bowl → Serve (with $m=4$), as shown in Figure 6.
+To extract subtask fractions and bottlenecks, we model each kitchen layout as a graph $G=(\mathcal{V},\mathcal{E})$, where nodes $\mathcal{V}$ represent spatial locations and workstations, and edges $\mathcal{E}$ represent valid movements.
+Each subtask $i$ corresponds to a path between workstations.
+We estimate each subtask fraction using normalized path lengths $f_{i}=d(i)/\sum_{j=1}^{m}d(j)$, where $d(i)$ is the length of the shortest path needed to complete subtask $i$.
+We assume a fixed cost of 10 steps for Cook.
+We then compute concurrency limits per subtask.
+The spatial bottleneck $C^{s}_{i}$ is based on edge betweenness centrality along subtask $i$’s path, reflecting the narrowest point in available throughput [40].
+The resource bottleneck $C^{r}_{i}$ is defined as the number of agents that can simultaneously use the workstation(s) required for subtask $i$.
+The overall concurrency limit is $C_{i}=\min(C^{s}_{i},C^{r}_{i})$.
+These values are substituted into Equation 3 to compute parallelizability $S(N,B)$.
+For full computation details, see Appendix D.
+4 Experiment 1: Divergent Specialization Patterns Between Environments
+Figure 2: Experiment 1 Results.
+(A) In SMAC, $N$ agents cooperate to defeat enemy units.
+(i) With no bottlenecks, parallelizability predicts $S(N,C)=N$.
+(ii) Specialization (SI) decreases with increasing parallelizability, as predicted (r=0.86r=-0.86).
+(B) In MPE, agents cooperate to cover $m$ landmarks that can be occupied by 1 agent.
+(C) Across environments, task parallelizability $S$ is strongly inversely correlated with specialization (r=0.84r=-0.84), and the logistic regression classifier achieves perfect accuracy ($\text{accuracy}=1.0$).
+In Experiment 1, we test whether task parallelizability can predict deviations between specialist and generalist policies in vastly different environments, providing empirical validation for our framework.
+In SMAC, where agents face minimal spatial and resource constraints, generalist policies prevail and specialization further decreases with team size, consistent with high task parallelizability.
+In contrast, MPE imposes strict bottlenecks since each landmark can only be occupied by a single agent, resulting in high specialization.
+These results show that parallelizability plays a critical role in shaping the emergence of specialist or generalist behavior.
+4.1 Experiment Design
+We evaluated teams of cooperative agents on two benchmark environments using JaxMARL [41].
+In both environments, agents were trained independently using a recurrent actor–critic network and Independent PPO (IPPO) [42, 43].
+Full training details are provided in Appendix E.
+In SMAC, we systematically varied team size ($N\in\{2,3,4,5,6\}$) and task complexity, defined as the number of enemy units per episode ($m\in\{2,3\}$).
+Each $(N,m)$ configuration was trained with 10 random seeds, resulting in 100 total runs.
+In MPE, we varied team size ($N\in{2,3,5,10}$) and task complexity, defined as the number of target landmarks ($m\in{2,3,5,10}$) which agents must cover simultaneously.
+Each $(N,m)$ configuration was trained with 10 random seeds, resulting in 160 total runs.
+4.2 Results
+Distribution of SI
+Results are depicted in Figure 2.
+Experiment 1 revealed a clear relationship between task parallelizability and specialization across environments.
+In SMAC, agents always converged to generalist policies (mean SI =0.06=0.06), with all trial-level mean SIs below $0.09$ and all raw SIs below $0.20$.
+Specialization decreased with increasing team size: the mean SI across agents was negatively correlated with $N$ (Pearson r=0.86r=-0.86).
+This aligns with the prediction that SMAC has no task bottlenecks ($S(N,C)=N$), allowing high parallelizability and low specialization.
+In contrast, MPE exhibited consistently high specialization across all conditions (mean SI =0.61=0.61).
+Even in the best-performing seeds, SI remained above $0.5$, consistent with the environment’s strict spatial bottlenecks.
+Because each of the landmarks must be uniquely covered by one agent, the number of subtasks equals the number of agents and no two agents can complete the same subtask.
+Our parallelizability metric thus assigns $S=1$ for MPE, predicting maximum specialization.
+Model Fit
+To test the predicted relationship, we pooled data across environments and used logistic regression to predict whether agents would specialize based on the parallelizability score $S$.
+This model achieved perfect classification accuracy (1.0) across all trials.
+The relationship between $S$ and specialization was also strongly linear, with a Pearson correlation of r=0.84r=-0.84 ($p=5.9\times 10^{-8}$).
+5 Experiment 2: Analysis of Mixed Bottlenecks in Overcooked-AI
+Figure 3: Experiment 2 Results.
+(A) Six $5\times 5$ layouts with varying bottlenecks were paired with different recipes and pots for a total of 144 trials, 1440 runs.
+(B) Histogram of specialization (SI) in the best-performing seeds reveals a distinctly bimodal distribution, with most teams exhibiting low (generalist) or high (specialist) SI.
+(C) Scatter plot of task parallelizability ($S$) vs.
+observed SI.
+The negative correlation confirms that greater parallelizability corresponds to lower specialization.
+Experiment 2 shows that both spatial and resource bottlenecks influence specialization, and that our parallelizability metric remains predictive under complex constraints.
+These findings suggest that parallelizability can be used to systematically design environments that promote or inhibit specialization, without the need to tailor training algorithms.
+When agents are given ample space and workstations, they develop generalist policies and learn the full scope of the task; in contrast, spatial or resource constraints lead to specialist policies.
+To test these predictions, we use Overcooked-AI, which features richer spatial topologies and resource constraints, supporting a wide range of strategies.
+5.1 Experiment Design
+We constructed six $5\times 5$ layouts with varying spatial bottlenecks: (1) open space with surrounding counters, (2) single vertical divider, (3) double vertical divider, (4) horizontal divider, (5) two dividers, and (6) long divider, illustrated in Figure 3.
+We varied resource bottlenecks (number of pots) while controlling for relative distances between workstations.
+Each layout featured four workstation permutations, totaling 48 configurations.
+Agents were trained to make one-, two-, or three-onion soup with 10 random seeds for a total of 1440 runs.
+For each trial, we trained two independent agents collaboratively via self-play using JaxMARL and iPPO [42, 43, 41].
+Training and compute details are in Appendix E.
+We selected the highest-performing seed based on joint reward during the roll-out.
+If multiple seeds reached the same reward, we averaged their SI.
+Trials were excluded if an agent failed to perform any successful action or the team scored a reward of 0.
+Four trials were excluded.
+5.2 Results
+Results are shown in Figure 3.
+Our simulations revealed a clear bimodal distribution of specialization.
+There were 81 specialist teams with $SI\geq 0.5$ ($57.86\%$) and 59 generalist teams ($42.14\%$).
+Our task parallelizability model was generally successful at predicting specialization.
+The Pearson correlation between predicted task parallelizability and SI was 0.67-0.67 ($p=8.94\times 10^{-6}$), affirming the relationship between parallelizability and specialization.
+To validate the model’s predictive power, we used logistic regression to classify high ($SI\geq 0.5$) or low specialization ($SI<0.5$).
+The model achieved a test set accuracy of $90.90\%$, with $0.88$ precision ($F1=0.93$) for high specialization and $1.00$ precision ($F1=0.86$) for low.
+Details are in Appendix F.
+6 Experiment 3: Exploratory Large-Scale Analysis in Overcooked-AI
+In Experiment 2, we manipulated spatial and resource bottlenecks in controlled Overcooked environments, holding layout size and workstation counts (except for pots) constant.
+To more thoroughly examine how task parallelizability drives specialization, Experiment 3 expands the analysis by systematically varying layout size, bottleneck severity, and the number and arrangement of workstations.
+6.1 Experiment Design
+Figure 4: Experiment 3 Results.
+(A) Histogram of SI in best-performing seeds for 3,200 Overcooked configurations confirms bimodal distribution of generalists and specialists.
+(B) Variables influencing parallelizability show strong, significant effects on SI: specialization decreases greater concurrency (more pots, fewer spatial bottlenecks) and increases with subtasks.
+Error bars represent SEM.
+We generated a test suite of 3,200 unique Overcooked environments for a large-scale exploratory analysis.
+We created 14 spatial configurations with unique arrangements of counters, including the original layouts from Overcooked-AI and a new set of layouts inspired by real-world kitchens [44].
+A complete set of layouts is provided in Figure 7.
+To create diverse parallelizability constraints, we systematically varied the number of workstations (onions, tomatoes, pots, bowls, and serving stations).
+A total of 16 combinations included one or two onion piles, bowls, pots, serving stations, and zero or one tomatoes.
+Workstation positions were randomized three times per layout.
+Agents were trained to make (i) 1-onion (ii) 2-onion (iii) 3-onion (iv) 1-onion and 1-tomato and (v) 2-onion and 1-tomato soup.
+These variations produced a 3,200 unique environments.
+We used the training procedures from Experiment 2, and again selected the highest-performing seed.
+6.2 Results
+Distribution of SI.
+Results again revealed a distinctly bimodal distribution, yet more skewed towards specialization.
+Specialists ($SI\geq 0.5$) dominated $74.9\%$ of runs, with $48.9\%$ fully specialized ($SI>0.9$).
+Generalists ($SI<0.5$) appeared in $25.1\%$ of runs, including $22.3\%$ fully generalist ($SI<0.1$).
+The remaining $28.8\%$ exhibited mixed specialist-generalist behaviors ($0.1\leq SI\leq 0.9$).
+Model Evaluation
+Statistics can be found in Appendix G and H, Tables 3 and 4.
+For each of the variables presented in our predictive model, there was a significant effect in the expected direction on observed SI.
+We found a significant effect of the number of pots, but not the total number of workstations; spatial bottlenecks; and recipe complexity (the number of ingredients required to make a soup) on emergent specialization.
+The average correlation 0.486-0.486 ($p=8.49\times 10^{-28}$) revealed a moderate negative relationship between parallelizability and SI.
+A logistic regression classifying low vs.
+high specialization ($SI\geq 0.5$) achieved $74.2\%$ accuracy, with $0.93$ precision ($F1=0.81$) for high specialization and $0.46$ precision ($F1=0.58$) for low specialization.
+Effects of State Space Size
+As illustrated, our model predictions deviated more from observed SI in this larger-scale analysis.
+Further investigation revealed a moderate correlation between grid size and SI ($r=0.383,p=5.57\times 10^{-17}$), even when controlling for bottlenecks (Figure 8).
+In larger layouts, agents tended to converge to specialist policies, even with high predicted parallelizability.
+7 Experiment 4: Deviations from MARL Training Biases
+Figure 5: Experiment 4 Results.
+Experiment 2 and parallelizability predictions reveal that generalist policies are most efficient in each layout.
+However, specialization significantly increases with layout size, also leading to lower reward for more complex tasks.
+In Experiment 4, we test whether deviations from our predictions in Experiment 3 stem from limits of our parallelizability framework or from increased exploration demands in larger state spaces.
+We find that as state space grows, despite a fixed grid size that is needed to complete the task, agents adopt more specialized strategies and achieve lower rewards with larger recipes.
+This suggests that in complex environments, specialization serves as a tractable alternative, enabling agents to focus on restricted subspaces when optimal joint policy learning is challenging.
+7.1 Experiment Design
+We re-used the first four $5\times 5$ layouts with two pots from Experiment 2 (depicted in Figure 5) and systematically varied seven layout sizes and two conditions.
+In the constant condition, we increased the overall grid size but held the amount of floor space that agents could access constant.
+In the non-constant condition, we increased the overall grid size and allowed agents to access it.
+Critically, Experiment 2 confirmed that the optimal behavior for these layouts were generalists; although the size of the state space differs, the path lengths required to actually complete the tasks remained the same.
+In total, there were 280 runs (7 sizes, 2 conditions, 2 recipes, and 10 random seeds).
+7.2 Results
+Distribution of SI
+For the one-onion recipe, agents consistently converged to fully specialist policies in three of the layouts ($100\%$ with a single divider, $81.3\%$ with a double divider, and $81.3\%$ with no-divider), whereas only $46.3\%$ of runs in the counter circuit resulted in SI $\geq 0.5$.
+For the three-onion soup recipe, agents also consistently converged to specialist policies ($93.8\%$ with a single divider, $87.5\%$ with a double divider, $87.5\%$ with no-divider, and $93.8\%$ in the counter circuit).
+Interaction Between State Space Size and Task Complexity
+The effect of layout size varied across the two recipes, indicating that both state space size and the number of tasks create a more difficult learning problem.
+With one-onion soup, SI increased significantly with layout size ($r=0.7924,p=0.03$).
+While there was a negative relationship between throughput and improper specialization, it was not statistically significant (r=0.2512,p=0.5869r=-0.2512,p=0.5869).
+In contrast, for the three-onion soup task, SI increased significantly with layout size ($r=0.7549,p=0.05$) and throughput significantly decreased (r=0.8556,p=0.0140r=-0.8556,p=0.0140), confirming that deviations from generalist strategies when possible lead to worse team performance (Proposition 4).
+8 Related Work
+The study of specialization in multi-agent systems spans multiple disciplines.
+Division of labor leading to functional specialization has been widely studied in biological and social domains, including cells [45, 46], insect colonies [47, 48], and human societies [49, 50, 14, 17].
+Inspired by these insights, research in MARL and robotics has sought to design algorithms that actively encourage specialization by promoting policy diversity [51, 4, 5, 6, 7, 8, 9, 10, 12] or by organizing agents into hierarchies [52, 53].
+Similarly, new work on language-agent collaboration often assumes that specialization is necessary, prompting agents to adopt distinct roles [54, 55, 13, 56, 11].
+However, while biological and social systems provide compelling examples, much of this work assumes that specialization is beneficial without systematically identifying when and why generalists might outperform specialists.
+For example, one benefit of generalists is their flexibility, which allows them to outperform rigidly specified roles under conditions of high environmental variability or individual failure [57, 16].
+A useful contrast comes from distributed systems, which are explicitly engineered rather than emergent.
+In these systems, heterogeneity and parallelism are carefully designed to optimize performance [58, 20, 59] and minimize contention [60].
+This structured approach provides a useful reference point for understanding when emergent specialization is not just incidental, but optimal.
+In our work, we propose a novel theoretical framework for predicting the utility of specialization based on task efficiency, considering environmental bottlenecks, resource constraints, and task complexity.
+Prior work in RL has shown that spatial bottlenecks such as doorways can shape subgoals [61], while research in human-robot collaboration and MARL has demonstrated that environmental design influences emergent behavior [62, 63].
+Building on these insights, we formalize a predictive model for when generalist versus specialist strategies should emerge.
+9 Discussion
+Specialization, or the differentiation of agents’ policies and roles, is often assumed to improve team performance.
+However, principles from distributed systems offer an alternative hypothesis.
+Specifically, generalist agents may outperform specialists when concurrent execution of the task would improve speed and efficiency.
+In this work, we outlined a theoretical framework for predicting generalist versus specialist efficiency in cooperative multi-agent tasks inspired by a variant of Amdahl’s Law.
+We then validated this model across core MARL environments (SMAC and MPE), as well as three in-depth Overcooked-AI experiments.
+Together, these findings suggest that generalists are constrained by their environment and task parallelizability, making it possible to design layouts which either encourage or inhibit role emergence regardless of whether the training algorithm explicitly incentivizes diversity.
+Our framework is highly scalable, providing principled predictions for emergent behavior given various team sizes, environments, and task constraints.
+More broadly, this framework motivates the use of distributed systems in analyzing trade-offs in multi-agent systems.
+This work has several limitations.
+First, our model does not account for switching behavior (e.g., generalists who alternate roles) or cases where specialist and generalist policies perform equally well, reflecting an implicit assumption from Amdahl’s Law that subtasks are executed sequentially without overlap.
+Second, while our model provides a theoretical upper bound on the benefits of generalists, it omits coordination overhead, which may make the empirical relationship non-monotonic.
+Third, our model assumes agents are initially homogeneous, yet specialization can also arise when agents have heterogeneous skills that make them more efficient at certain tasks [58, 64].
+Finally, Experiment 4 suggests that specialization is a locally optimal solution that agents use when exploration is costly, i.e.
+in larger state spaces.
+Clarifying this trade-off can provide insight into how specialization may improve sample efficiency, as well as how to decrease coordination overhead between agents.
+Appendix A State-Visitation and Action Distributions
+In order to measure the differences between agents’ policies, we must quantify how different their distributions of states and actions are.
+Each agent follows a state-visitation distribution that describes how frequently they occupy different states under their policy [26] given by:
 $d^{\pi}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s|\pi)$ (4)
@@
 This distribution captures both what actions an agent tends to take and how often they visit states where those actions are available.
-We approximate specialization using the Jensen-Shannon divergence (JSD) over agents’ action distributions $P_{i}$ (Endres & Schindelin, 2003; Fuglede & Topsoe, 2004).
-Given $N$ agents with action distributions $P_{1},P_{2},...,P_{N}$, JSD is given by:
-$\text{JSD}(P_{1},P_{2},...,P_{N})=\sum_{i=1}^{N}w_{i}D_{\mathrm{KL}}(P_{i}||M),$ (4)
-where $M=\sum_{i=1}^{N}w_{i}P_{i}$ is the mixture distribution, $w_{i}=\frac{1}{N}$, and $D_{\mathrm{KL}}(P_{i}||M)$ is the Kullback-Leibler (KL) divergence between each agent’s action distribution $P_{i}$ and $M$.
-When $\text{JSD}(P_{1},P_{2},...,P_{N})=0.0$, all $N$ agents have performed the same proportions of actions, indicating that all agents have performed all subtasks equally (no specialization).
-When $\text{JSD}(P_{1},P_{2},...,P_{N})=1.0$, the action distributions are completely disjoint (i.e., each agent assigns probability exclusively to a unique set of actions with no overlap), meaning that the agents specialize in complementary subtasks.
-Intermediate values of JSD indicate partial specialization, in which the agents overlap in some subtasks but not others.
-We use Jensen-Shannon rather than Kullback-Leibler divergence because it is symmetric, i.e.
-$\text{JSD}(P_{1}||P_{2})=\text{JSD}(P_{2}||P_{1})$.
-The degree of specialization in each team should be the same regardless of which agent specializes in which subtasks.
-2.3 Graphical Representations of Spatial Layouts
-To model task execution in structured environments, we represent spatial layouts as graphs.
-This allows for efficient computation of movement paths, bottlenecks, and spatial constraints on multi-agent coordination.
-We define a layout graph $G=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of walkable positions and task-relevant objects and $\mathcal{E}$ is the set of valid movements between adjacent positions.
-This spatial representation applies broadly to spatio-temporal multi-agent environments, including gridworlds.
-$G$ defines the static structure of the environment and constrains the MDP, where the state space $\mathcal{S}$ includes agent positions $x\subseteq\mathcal{V}$ and task-relevant object states.
-The transition function $\mathcal{T}$ ensures that agents execute valid movements along edges in $\mathcal{E}$.
-Definition 2.1.
+Appendix B Original Form of Amdahl’s Law
+Let $f$ represent the proportion of a task that can be parallelized and accelerated by a factor of $s$, while 1f1-f denotes the portion that must be executed serially.
+The theoretical speedup $S$ achievable when distributing a task across multiple processors is given by:
+$S=\frac{1}{(1-f)+\frac{f}{s}}$ (7)
+Appendix C Properties
+We can now examine several useful properties of this speed-up bound $S(N,C)$.
+These propositions formalize the intuition that removing bottlenecks promotes generalists, never specialists, since only generalists benefit from concurrent access to subtasks.
+Proposition 1 (Monotonicity). $S(N,B)$ is non-decreasing in each $C_{i}$.
+Proof. $s_{i}(N,C_{i})$ is non‑decreasing in $C_{i}$ and appears in the denominator of $S$; increasing any $C_{i}$ weakly decreases the denominator and therefore weakly increases $S$.
+Proposition 2 (Full concurrency favors generalists).
+Assume negligible task–switch costs.
+If $min_{i}C_{i}\geq N$ (every subtask can accommodate all agents), then $S(N,B)=N$ and any throughput‑maximizing policy set satisfies $\mathrm{SI}=0$.
+Proof.
+When $C_{i}\geq N$ for all $i$, we have $s_{i}(N,C_{i})=N$.
+Substituting into the definition of $S$ gives us $S(N,B)=1/\!\sum_{i}f_{i}/N=N$.
+Because all agents can execute every subtask concurrently, restricting any agent’s action set cannot improve throughput; the optimum is therefore achieved by identical (generalist) policies, for which $\mathrm{SI}=0$.
+Remark (Switching costs).
+When subtask times vary across workstations or switching between subtasks incurs large delays, the guarantee for generalist policies no longer holds; specialists may achieve higher throughput.
+Proposition 3 (Bottlenecks induce specialization).
+If $S(N,B)<N$, then no fully generalist team can attain throughput proportional to $N$; a role allocation with $\mathrm{SI}>0$ is required for optimality.
+Proof.
+$S<N$ implies $\exists\,i$ with $C_{i}<N$.
+Under identical generalist policies, the expected number of agents attempting subtask $i$ at any instant is $N$, but at most $C_{i}$ can make progress; the remainder are idle.
+This idle fraction reduces realized speed‑up below $N$.
+A specialist allocation that assigns at most $C_{i}$ agents to subtask $i$ and redistributes the rest to other subtasks eliminates that idle time.
+Hence any throughput‑optimal policy set must have $\mathrm{SI}>0$.
+Figure 6: Example of a task graph for making one-onion soup in Overcooked. Task-relevant objects are represented as nodes in a directed acyclic graph (DAG), and edges represent subtasks.
+Appendix D Layout Graphs
 A layout graph is an unweighted, directed graph $G=(\mathcal{V},\mathcal{E})$ where:
-$V=\{u\mid u\in\{(i,j)\mid i\in[1,n],j\in[1,m]\}\setminus\text{obstacles}\}$
+$\mathcal{V}=\{u\mid u\in\{(i,j)\mid i\in[1,n],j\in[1,m]\}\setminus\text{obstacles}\}$
 An edge $(u,v)\in\mathcal{E}$ exists if and only if $u$ and $v$ are adjacent, i.e., they share a horizontal or vertical edge in the grid.
@@
 (ii) If $v$ is a task-relevant object (e.g., workstation) there is a unidirectional edge $(u,v)\in\mathcal{E}$, but $(v,u)\notin\mathcal{E}$, reflecting the fact that agents can interact with but not move through it.
-2.4 Task Decomposition as a Directed Acyclic Graph
-To analyze task parallelizability, we represent task execution as a directed acyclic graph (DAG), where nodes denote subtasks and edges represent the order in which subtasks must be completed (Solway et al., 2014).
-Definition 2.2.
-We define a task graph $T=\{\mathcal{B}_{r},\mathcal{E}_{T}\}$, where $\mathcal{B}_{r}$ (subtask nodes) represent intermediate steps such as workstations or resources, and $\mathcal{E}_{T}$ indicate ordering constraints between subtasks (e.g., an agent must retrieve an item before assembling it).
-The number of subtasks $m$ depends on task complexity.
-Using the layout graph $G$, we determine the spatial locations and paths between subtasks.
-The fraction $f(i)$ of the total task spent on a given subtask $i$ is:
-$f(i)=\frac{d(i)}{\sum_{j=1}^{m}d(j)},$ (5)
-where $d(i)$ is the path length required to complete subtask $i$.
-2.5 Spatial and Resource Bottlenecks
-We introduce a framework to identify bottlenecks that constrain parallelizability in multi-agent tasks.
-This idea is inspired by sequential dependencies, resource contention, and communication overhead that prevent scalability in parallel processing (Almasi & Gottlieb, 1994; McCool et al., 2012; Zhuravlev et al., 2010).
-A classic example is memory bandwidth in GPUs, where increasing the number of threads does not lead to proportional speed-up due to limited shared access (Dublish et al., 2017).
-We extend this framework to multi-agent tasks by identifying two key types of bottlenecks: (1) spatial bottlenecks ($\mathcal{B}_{s}$) limit movement of multiple agents due to obstacles, narrow passages, and congestion, and (2) resource bottlenecks ($\mathcal{B}_{r}$) limit access to critical resources such as workstations or tools needed to complete subtasks (Mieczkowski et al., 2024).
-Together, these bottlenecks determine how effectively agents can execute tasks in parallel.
-They arise across numerous multi-agent domains, including warehouse operations (insufficient space, storage locations, and pallets (Živičnjak et al., 2022)), autonomous vehicles (lane reductions, traffic congestion (Zhang & Gao, 2020)), and smart factories (Wang et al., 2016).
-We represent spatial and resource bottlenecks as two global sets over $G=(\mathcal{V},\mathcal{E})$: $\mathcal{B}_{s}\subseteq\mathcal{E}$, $\mathcal{B}_{r}\subseteq\mathcal{V}$, where $\mathcal{B}_{s}$ consists of edges in $\mathcal{E}$ that constrain movement and $\mathcal{B}_{r}$ consists of nodes in $\mathcal{V}$ with task-relevant objects (e.g., workstations).
-Each subtask $i$ is constrained by a subset of bottlenecks relevant to the path ($\mathcal{B}_{s}(i)$) and resources ($\mathcal{B}_{r}(i)$) required to complete it.
-To quantify spatial bottlenecks, we compute edge betweenness centrality, a graph-theoretic measure of congestion where high-centrality edges indicate critical bottlenecks in connectivity (Barrat et al., 2004).
-Definition 2.3.
+We represent spatial and resource bottlenecks as two global sets over $G=(\mathcal{V},\mathcal{E})$: $C^{s}\subseteq\mathcal{E}$, $C^{r}\subseteq\mathcal{V}$, where $C^{s}$ consists of edges in $\mathcal{E}$ that constrain movement and $C^{r}$ consists of nodes in $\mathcal{V}$ with task-relevant objects (e.g., workstations).
+Each subtask $i$ is constrained by a subset of bottlenecks relevant to the path ($C^{s}_{i}$) and resources ($C^{r}_{i}$) required to complete it.
+To quantify spatial bottlenecks, we compute edge betweenness centrality, a graph-theoretic measure of congestion where high-centrality edges indicate critical bottlenecks in connectivity [40].
 The betweenness centrality of an edge $e$ in graph $G=(\mathcal{V},\mathcal{E})$ is:
-$C_{B}(e)=\sum_{s\neq t\in V}\frac{\phi(s,t|e)}{\phi(s,t)},$ (6)
+$B(e)=\sum_{s\neq t\in V}\frac{\phi(s,t|e)}{\phi(s,t)},$ (8)
 where $\phi(s,t)$ is the number of shortest paths between nodes $s$ and $t$, and $\phi(s,t|e)$ is the number of those paths that pass through edge $e$.
-Since not all bottlenecks affect every subtask, we compute a per-subtask spatial bottleneck score by summing the centralities of all edges along the shortest path required for that subtask: $\mathcal{B}_{s}(i)=\sum_{e\in\text{path}(u,v)}C_{B}(e)$, where path$(u,v)$ is the shortest path between subtask-relevant nodes $u$ and $v$.
-A higher $\mathcal{B}_{s}(i)$ indicates stronger spatial constraints that may inhibit parallel agent movement.
+Since not all bottlenecks affect every subtask, we compute a per-subtask spatial bottleneck score by summing the centralities of all edges along the shortest path required for that subtask: $C^{s}_{i}=\sum_{e\in\text{path}(u,v)}B(e)$, where path(u,v)(u,v) is the shortest path between subtask-relevant nodes $u$ and $v$.
+A higher $C^{s}_{i}$ indicates stronger spatial constraints that may inhibit parallel agent movement.
 To quantify resource bottlenecks, we simply sum the number of task-relevant nodes that are required for a subtask $i$.
 For example, in a manufacturing system there may be only two welding stations; thus, only two agents may complete that subtask at the same time.
-Formally, for each subtask $i$, we define a resource bottleneck score $\mathcal{B}_{r}(i)$ as: $\mathcal{B}_{r}(i)=\sum_{v\in\mathcal{B}_{r(i)}}c(v)$, where $c(v)$ is the capacity of each task-relevant node (or the maximum agents that can use it at once).
-2.6 Task Parallelizability Prediction
-We now adapt Amdahl’s Law, a measure of the expected speed-up when dividing a task across multiple processors given the proportion that must be executed serially, to multi-agent tasks (Amdahl, 1967; Hill & Marty, 2008; Cassidy & Andreou, 2011).
-For the original form of Amdahl’s Law, see Appendix A.1.
-Mieczkowski et al.
-(2024) extended this idea to predict idleness during human collaborations given the current group size, workload, and resource constraints.
-We hypothesize that a similar formulation can be used to predict specialization in multi-agent tasks.
-Let $f(i)$ represent the fraction of the total task time required for the $i^{th}$ of $m$ subtasks such that $\sum_{i=1}^{m}f(i)=1$.
-For each subtask $i$, the overall bottleneck capacity is defined as the minimum bottleneck that prevents agents executing it concurrently, or $\mathcal{B}_{i}=\min(\mathcal{B}_{s}(i),\mathcal{B}_{r}(i))$.
-The subtask speed-up factor $s_{i}(N,\mathcal{B}_{i})$ quantifies the efficiency gain from adding an $N$th agent to a subtask, given that $N-1$ agents are already working on it and $\mathcal{B}_{i}$ bottlenecks restrict parallel execution.
-Formally:
-$s(N,\mathcal{B}_{i})=\frac{\text{min}(N,\mathcal{B}_{i})}{\text{min}(N-1,\mathcal{B}_{i})}$ (7)
-Thus, if we assume that there is no cost to switching between subtasks, the overall expected speed-up achievable when an additional agent performs the entire task in parallel, which we call task parallelizability, is expressed as:
-$S(N,\mathcal{B})=\frac{1}{\sum_{i=1}^{m}\frac{f(i)}{s(N,\mathcal{B}_{i})}}$ (8)
-2.7 Task Parallelizability and Specialization
-Our task parallelizability prediction estimates the efficiency gain if an additional agent independently completed all $m$ subtasks in parallel.
-We propose that this measure sets a fundamental upper bound on the benefits of specialization in multi-agent systems.
-If $S(N,\mathcal{B})$ is low due to resource or spatial bottlenecks, then the task would not be significantly sped up by multiple agents executing it in parallel.
-In this case, there would be little value to a team of generalist agents, so if we want to increase task efficiency it is better for agents to specialize and perform complementary subtasks.
-However, specialization itself can introduce new bottlenecks, particularly when agents must coordinate frequently or wait for specific subtasks to be completed before proceeding.
-In environments with high parallelizability $S(N,\mathcal{B})$, generalist agents who work independently may avoid these delays, leading to faster performance.
-3 Application to Overcooked-AI
-To investigate how task parallelizability influences specialization, we apply our theoretical model for predicting generalist efficiency to a specific multi-agent environment (Carroll et al., 2019).
-Overcooked-AI is a fully cooperative two-agent platform in which pairs of agents must deliver soups as quickly as possible while navigating obstacles, avoiding collisions, and coordinating subtasks (Wu et al., 2021).
-Each recipe requires a different number of subtasks ($m$) to prepare a soup, while resource bottlenecks such as the number of pots ($\mathcal{B}_{r}$) and spatial bottlenecks such as narrow passages ($\mathcal{B}_{s}$) constrain task parallelizability.
-Layout Graphs. We represent each layout graph according to Definition 2.1. An example is depicted in Figure 2.
-Figure 2: Examples of Overcooked layouts, layout graphs, and spatial bottleneck scores $\mathcal{B}_{s}(i)$ for subtask Onion $\to$ Pot.
-Layout A is an open room with low spatial bottlenecks, whereas B has a narrow passage leading to the pot, creating a greater bottleneck.
-Task Decomposition.
-We decompose the overall task of “cooking a soup” using a directed acyclic graph (DAG) $T=(\mathcal{B}_{r},\mathcal{E}_{T})$ according to Definition 2.2, where $\mathcal{B}_{r}$ is the set of resources, or workstations (e.g., onion piles, pots, bowls), and $(m_{i}\to m_{j})\in\mathcal{E}_{T}$ is a directed edge representing a required subtask (such as moving an onion to the pot).
-An example can be found in Figure 5.
-The number of subtasks $m$ depends on the recipe.
-To make a one-onion soup, there are $m=4$ subtasks total: Onion $\to$ Pot, Cook in Pot, Bowl $\to$ Pot, and Pot $\to$ Serve.
-To make a two-onion or one-onion one-tomato soup, there are $m=5$ subtasks.
-To make a three-onion or two-onion one-tomato soup, there are $m=6$ subtasks.
-Using the layout graph $G$, we determine the locations and paths between workstations.
-The fraction $f(i)$ spent on subtask $i$ is its path length relative to the total path length of all subtasks following Equation 5.
-Bottlenecks.
-We manipulate and measure bottlenecks in Overcoooked in two ways.
-We vary resource bottlenecks $\mathcal{B}_{r}$ by changing the number of workstations ($p$ pots, $o$ onions, $t$ tomatoes, $d$ bowls, and $s$ serving stations) available to agents.
-We vary spatial bottlenecks $\mathcal{B}_{s}$ by modifying the layout size and arrangement of obstacles (represented as counters that agents cannot cross).
-We quantify these spatial bottlenecks for each subtask $i$ using edge betweenness centrality as described in Definition 2.3.
-Specialization.
-To measure specialization, we convert agents’ behaviors (move, interact, wait) from a policy roll-out of the best-performing seed to high-level actions: (i) picking up onions, (ii) placing onions in pot, (iii) picking up tomatoes, (iv) placing tomatoes in pot, (v) picking up bowls, (vi) moving soup from pot to bowl, and (vii) serving soup.
-Each agent $i\in\mathcal{N}$ has an action distribution $P_{i}:\mathcal{A}\to[0,1]$, where $P_{i}(a)$ represents the probability of agent $i$ taking action $a\in\mathcal{A}$ during a policy roll-out after training.
-For each agent $i$, we compute their high-level action distribution $P_{i}$ by normalizing the frequency of each action over the total actions they performed during the roll-out.
-We measure specialization as the Jensen-Shannon divergence between $|\mathcal{N}|=2$ agents, $\text{JSD}(P_{1}||P_{2})$.
-Task Parallelizability Prediction.
-Our model for predicting task parallelizability measures the relative efficiency gained from two generalist agents making soups in parallel versus specializing in subtasks (e.g., fetching ingredients versus serving soup).
-Consider the task of making one-onion soup.
-This requires $m=4$ subtasks: Onion $\to$ Pot, Cook in Pot, Bowl $\to$ Pot, and Pot $\to$ Serve.
-Each subtask $i$ is associated with (i) a number of workstations $\mathcal{B}_{r}(i)\in\mathcal{B}_{r}$, (ii) a path length $d(i)$, and (iii) a measure of spatial bottlenecks $\mathcal{B}_{s}(i)$ between those workstations, where $\mathcal{B}_{s}(i)\geq 0.8$ indicates constrained areas.
-If there are multiple paths, we consider $d_{\min}(i)$.
-Subtask durations are weighted by path length (Equation 5), with Cook assumed to take 10 time-steps.
-The overall speed-up expected from two generalist agents making soups in parallel is thus given by $S(N,\mathcal{B})$ in Equation 8.
-We hypothesize an inverse relationship between $S(N,\mathcal{B})$ and $\text{JSD}(P_{1}||P_{2})$.
-This would suggest agents specialize more when task and environmental constraints inhibit them from executing the same subtasks simultaneously.
-4 Experiment 1: Exploring Specialization in Large-Scale Overcooked Simulations
-Figure 3: Experiment 1 Results.
-(A) Histogram of specialization (JSD) in the best-performing seeds for 3,200 unique Overcooked configurations reveals a bimodal distribution, with teams clustering as either fully generalist or specialist.
-(B) Scatter plot of predicted task parallelizability ($S$) vs.
-observed JSD reveals a moderate negative correlation, indicating less specialization when a task has greater parallelizability.
-(C) Variables influencing parallelizability show strong, significant effects on JSD: specialization decreases with more pots but increases with greater spatial bottlenecks and task complexity.
-Error bars represent standard error.
-In Experiment 1, we conduct a large-scale analysis across a variety of environments to test whether parallelizability impacts the emergence of specialization.
-4.1 Experiment Design
-We generated a test suite of 3,200 unique Overcooked environments for exploratory analysis.
-We utilized 14 spatial configurations representing 2D gridworld kitchens, each with a unique arrangement of counters and floor areas, including the original layouts from Overcooked-AI and a new set of layouts inspired by real-world kitchens (Pejic et al., 2019).
-A complete set of layouts is provided in Figure 6.
-To create diverse parallelizability constraints, we systematically varied the number of workstations (onions, tomatoes, pots, bowls, and serving stations).
-A total of 16 combinations included one or two onion piles, bowls, pots, serving stations, and zero or one tomatoes.
-Workstation positions were randomized three times per layout.
-Agents were trained to make five recipes: (i) 1-onion (ii) 2-onion (iii) 3-onion (iv) 1-onion and 1-tomato and (v) 2-onion and 1-tomato soup.
-These variations produced a total of 3,200 unique environments.
-For each of these environments, we trained two independent agents to play Overcooked collaboratively via self-play using PureJaxRL and PPO with five random seeds (Schulman et al., 2017; Lu et al., 2022).
-Full details regarding training procedures can be found in Appendix A.3.
-We then selected the highest-performing seed based on the joint reward that agents achieved during the roll-out.
-If multiple seeds reached the same maximum reward, we averaged their JSD.
-Trials were excluded if agents failed to take high-level actions or achieved a roll-out reward of 0.
-4.2 Results
-Distribution of JSD.
-Results from Experiment 1 are depicted in Figure 3.
-Our simulations revealed two distinct behavioral modes.
-Specialists ($JSD\geq 0.5$) dominated $74.9\%$ of runs, with $48.9\%$ fully specialized ($JSD>0.9$).
-Generalists ($JSD<0.5$) appeared in $25.1\%$ of runs, including $22.3\%$ fully generalist ($JSD<0.1$).
-The remaining $28.8\%$ exhibited mixed specialist-generalist behaviors ($0.1\leq JSD\leq 0.9$).
-Effects of Parallelizability on Specialization.
-For each of the variables presented in our predictive model, there was a significant effect in the expected direction on observed JSD.
-Full statistical details can be found in Appendix A.4.
-We found a significant effect of the number of pots, but not the total number of workstations; spatial bottlenecks; and recipe complexity (the number of ingredients required to make a soup) on emergent specialization.
-Model Evaluation.
-We analyzed the relationship between predicted task parallelizability and observed specialization (JSD) across 14 Overcooked layouts.
-Full statistical details are in Appendix A.5, with Tables 1 and 2 summarizing correlations, JSD, and rewards.
-The average correlation $-0.486$ ($p=8.49\times 10^{-28}$) revealed a moderate negative relationship between generalist efficiency and JSD, supporting our hypothesis that greater task parallelizability predicts lower specialization.
-A logistic regression classifying low vs.
-high specialization ($JSD\geq 0.5$) achieved $74.2\%$ accuracy, with $0.93$ precision ($F1=0.81$) for high specialization and $0.46$ precision ($F1=0.58$) for low specialization.
-Effects of State Space on Specialization We also observed a moderate correlation between grid size and average JSD ($r=0.383,p=5.57\times 10^{-17}$), even when controlling for resource and spatial bottlenecks (Figure 7).
-This suggests that state space size and workstation distances may act as confounding variables, influencing specialization independently of task-specific constraints.
-Larger state spaces increase exploration demands.
-In response, agents may adopt specialization as a locally optimal strategy to reduce policy complexity by focusing on specific paths and workstation interactions.
-5 Experiment 2: Validating Predictive Model
-Figure 4: Experiment 2 Results.
-(A) Six layouts (four depicted) with varying bottlenecks were designed on a $5\times 5$ grid, each with one or two pots, two onions, one bowl, and one serving station across four locations, yielding 48 layouts.
-Each layout was paired with one-, two-, or three-onion soup recipes for a total of 144 trials.
-(B) Histogram of specialization (JSD) across the best-performing seeds reveals a distinctly bimodal distribution, with most clustering around low (generalist) or high (specialist) JSD.
-(C) Scatter plot of predicted task parallelizability ($S$) vs.
-observed JSD averaged.
-A strong negative correlation confirms that greater potential speed-up from generalists corresponds to lower specialization.
-In Experiment 1, we confirmed that the variables which should affect generalist strategies from speeding up task completion have a significant impact on specialization.
-In Experiment 2, we validate our results using a smaller, controlled experiment.
-We also show that we can use our findings to design environments which will induce or inhibit the emergence of specialized roles.
-5.1 Experiment Design
-We utilized the same methods, game dynamics, and training schemes as in Experiment 1 (described in Section 4.1).
-To control state space size while varying spatial bottlenecks, we constructed six $5\times 5$ grid layouts: (1) open space with surrounding counters, (2) single vertical divider, (3) double vertical divider and (4) horizontal divider.
-These spatial layouts are illustrated in Figure 4.
-We varied the number of pots (1 or 2) rather than workstations, since pots had the largest effect on specialization in Experiment 1, while controlling for relative distances between workstations across trials.
-Each layout featured four workstation arrangements, totaling 48 configurations.
-For each, agents were trained to make one-, two-, or three-onion soup with 10 random seeds for a total of 1400 runs.
-As in Experiment 1, we again selected the highest-performing seed in terms of the joint reward that agents achieved during the policy roll-out after averaging over the randomized locations of workstations.
-Four trials were excluded from Experiment 2.
-5.2 Results
-Results from Experiment 2 are depicted in Figure 4.
-By fixing the size of the grid, the distinction between generalist and specialist strategies became more pronounced, with a clear bimodal distribution emerging.
-There were 81 specialist teams with $JSD\geq 0.5$ ($57.86\%$) and 59 generalist teams ($42.14\%$).
-Notably, the proportion of generalists increased substantially compared to Experiment 1 once we removed the confound of altering layout size, providing further evidence that specialization is partially driven by the size of the state space.
-Additionally, our predictive model achieved even greater accuracy for predicting specialization.
-The Pearson correlation between predicted task parallelizability and JSD was $-0.667$ ($p=8.94\times 10^{-6}$), confirming the hypothesized inverse relationship between generalist efficiency and specialization.
-Statistical details can be found in Appendix A.6.
-To validate the model’s predictive power, we fit a class-balanced logistic regression to classify high ($JSD\geq 0.5$) or low specialization ($JSD<0.5$).
-The model achieved a test set accuracy of $90.90\%$, demonstrating strong performance with $0.88$ precision ($F1=0.93$) for high specialization and $1.00$ precision ($F1=0.86$) for low.
-These results suggest that spatial bottlenecks and resource availability are key factors driving the emergence of specialization.
-Our findings indicate that our predictive model can be used to systematically design environments that will either promote or inhibit specialization.
-Rather than tailoring training algorithms to induce policy diversity, it is possible to do so based on environmental design alone.
-If agents are given ample space and workstations to learn the task, they will develop emergent generalist behaviors.
-Alternatively, if the environment is constrained spatially or in terms of resources, agents will develop specialist roles.
-6 Experiment 3: Investigating State Space Size and Specialization
-In Experiment 3, we investigate the relationship between state space size and specialization in greater detail.
-Our findings show that as the state space expands, even if the size of the grid that the agents can explore is held constant, agents consistently converge on more specialized strategies, indicating that specialization is a less computationally costly alternative when learning is made more difficult.
-These findings highlight a challenge in MARL: as state spaces grow, agents struggle to converge on optimal generalist policies.
-Full details, including the experimental design, results and statistical analysis are provided in Appendix A.7 and A.8.
-7 Related Work
-The study of specialization in multi-agent systems spans multiple disciplines.
-Division of labor leading to functional specialization has been widely studied in biological and social domains, including cells (Bell & Mooers, 1997; Rueffler et al., 2012), insect colonies (Ratnieks & Anderson, 1999; Fjerdingstad & Crozier, 2006), and human societies (Kitcher, 1990; Kreider et al., 2022; Goldstone et al., 2024; Vélez et al., 2024).
-Inspired by these insights, research in MARL and robotics has sought to design algorithms that actively encourage specialization by promoting policy diversity (Murciano et al., 1997; Padgham & Winikoff, 2002; Zhu & Zhou, 2008; Charakorn et al., 2020; Wang et al., 2020a, b; Li et al., 2021; Bettini et al., 2023, 2024) or by organizing agents into hierarchies (Zhang et al., 2010; Ahilan & Dayan, 2019).
-Similarly, new work on language-agent collaboration often assumes that specialization is necessary, prompting agents to adopt distinct roles before collaborating (Wang et al., 2023; Suzgun & Kalai, 2024; Swanson et al., 2024; Liu et al., 2024; Juang et al., 2024).
-However, much of this work assumes that specialization is always beneficial, without systematically identifying when and why generalist strategies may outperform specialists.
-A useful contrast comes from distributed systems, which are explicitly engineered rather than emergent.
-In these systems, heterogeneity and parallelism are carefully designed to optimize performance (Moncrieff et al., 1996; Almasi & Gottlieb, 1994; Pancake, 1996) and minimize contention (Park & Reveliotis, 2001).
-This structured approach provides a useful reference point for understanding when emergent specialization is not just incidental, but optimal.
-While biological and social systems provide compelling examples, it remains an open question when specialization emerges as the optimal architecture.
-Prior work suggests that under conditions of high environmental variability or individual failure, flexible generalist strategies may outperform rigidly specialized roles (Staps & Tarnita, 2022; Ben-Oren et al., 2023).
-In our work, we propose a novel theoretical framework for predicting the utility of specialization based on task efficiency, considering environmental bottlenecks, resource constraints, and task complexity.
-Prior work in RL has shown that spatial bottlenecks such as doorways can shape subgoals (Menache et al., 2002), while research in human-robot collaboration and MARL has demonstrated that environmental design influences emergent behavior (Fontaine et al., 2021; McKee et al., 2022).
-Building on these insights, we formalize a predictive model for when generalist versus specialist strategies should emerge.
-8 Discussion
-Specialization, or the differentiation of agents’ policies and roles, is often assumed to improve collaborative performance.
-However, principles from distributed systems offer an alternative hypothesis.
-Specifically, generalist agents may outperform specialist ones when concurrent execution of the entire task would improve speed and efficiency.
-In this work, we outlined a theoretical framework for predicting generalist versus specialist efficiency in cooperative multi-agent tasks inspired by a variant of Amdahl’s Law.
-We then validated this model across three Overcooked experiments, finding highly significant effects of parallelizability factors on emergent specialization and high predictive accuracy.
-Together, these findings suggest that specialization is constrained by the environment and parallelizability of a task, making it possible to design layouts which either encourage or inhibit role emergence regardless of whether the training algorithm explicitly incentivizes diversity.
-More broadly, this framework motivates the use of distributed systems in analyzing trade-offs that arise when dividing labor.
-This work has several limitations.
-First, our model does not account for switching behavior (generalists who switch roles rather than cooking independently) or cases where specialist and generalist policies are equally optimal.
-This limitation arises from an implicit assumption from Amdahl’s Law that subtasks are executed sequentially without overlap.
-Second, our model provides a theoretical upper bound on the benefits of generalists, but does not account for potential coordination overhead when increasing the number of agents, which could make the empirical relationship non-monotonic.
-Third, our model assumes agents are equally capable, yet specialization can also arise when agents have heterogeneous skills that make them more efficient at certain tasks (Moncrieff et al., 1996; Ginting et al., 2022).
-Despite these limitations, our predictive model is highly extensible.
-It generalizes from 2 to $n$ agents, providing explicit predictions for larger teams.
-Finally, our work suggests that specialization is a locally optimal solution that agents use when exploration is costly, i.e.
-in larger state spaces.
-Clarifying this trade-off can provide insight into how specialization may improve sample efficiency, as well as how to decrease coordination overhead between agents.
-Acknowledgments
-This work was supported in part by the Department of Defense (DoD) through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program, the Office of Naval Research (ONR) MURI grant (N00014-22-1-2813), the EPSRC CDT in RAS (EP/L016834/1), and a grant from the Templeton World Charity Foundation.
-Appendix A Appendix
-A.1 Original Form of Amdahl’s Law
-Let $f$ represent the proportion of a task that can be parallelized and accelerated by a factor of $s$, while $1-f$ denotes the portion that must be executed serially.
-The theoretical speedup $S$ achievable when distributing a task across multiple processors is given by:
-$S=\frac{1}{(1-f)+\frac{f}{s}}$ (9)
-A.2 Task Decomposition Graph
-Figure 5: Example of a task graph for making one-onion soup in Overcooked. Task-relevant objects are represented as nodes in a directed acyclic graph (DAG), and edges represent subtasks.
-A.3 Training Procedures
-For each trial in our MARL experiments, we trained two independent agents to play Overcooked collaboratively through self-play.
-Our implementation builds upon the PureJaxRL framework for Proximal Policy Optimization (Schulman et al., 2017; Lu et al., 2022).
-Each agent was implemented using an soft Actor-Critic network (Haarnoja et al., 2018), in which both the actor and critic networks were composed of two fully-connected layers with 64 hidden units each.
-We trained each agent using Proximal Policy Optimization with an Adam optimizer and a learning rate of $2.5\times 10^{-3}$ for $1\times 10^{7}$ timesteps (Zhang, 2018).
-At each timestep, the agents observed the state of the current environment as a $\text{width}\times\text{height}\times 27$-dimensional grid.
+Formally, for each subtask $i$, we define a resource bottleneck score $C^{r}_{i}$ as: $C^{r}_{i}=\sum_{v\in C^{r}_{i}}c(v)$, where $c(v)$ is the capacity of each task-relevant node (or the maximum agents that can use it at once).
+Appendix E Training Procedures
+For each trial, we trained two independent agents to play SMAC, MPE, or Overcooked collaboratively through self-play.
+Our implementation builds upon the PureJaxRL and JaxMARL frameworks for Proximal Policy Optimization [42, 43, 41].
+Each SMAC and MPE agent was implemented using a recurrent actor-critic network.
+Observations were embedded using a fully connected ReLU layer, followed by a gated recurrent unit (GRU) to model temporal dependencies.
+Each Overcooked-AI agent was implemented using an soft actor-critic network [65], in which both the actor and critic networks were composed of two fully-connected layers with 64 hidden units each.
+We trained each agent using Independent PPO with an Adam optimizer and a learning rate of $2.5\times 10^{-3}$ for $1\times 10^{7}$ timesteps [66].
+For SMAC, at each timestep, each agent received an observation vector containing local features about itself and other units within its field of view.
+This included its own health, position, unit type, and weapon cooldown, as well as relative information about nearby allies and enemies (e.g., health, positions, last movement direction, and attack status).
+The global state included information about all units on the battlefield and was only used during training.
+The policy produced a categorical distribution over a discrete set of actions, including movement, attack actions, and wait.
+For MPE, at each timestep, each agent received a continuous observation vector that encoded its own velocity and position, the relative positions of other agents, and the relative positions of static landmarks.
+Agents operated in a continuous 2D space.
+The policy output a categorical distribution over these actions, with masking used when necessary (e.g., to prevent collisions or enforce environment constraints).
+For Overcooked-AI, at each timestep, the agents observed the state of the current environment as a $\text{width}\times\text{height}\times 27$-dimensional grid.
 The first two dimensions corresponded to the spatial coordinates of the grid, while the 27-dimensional vector encoded the state of each grid cell, indicating whether it contained an agent, onion, tomato, bowl, or workstation.
 The policy produced a categorical distribution over discrete actions.
-To accelerate learning in the sparse-reward environments, we applied reward shaping to augment raw rewards with an auxiliary-shaped signal.
+To accelerate learning in the sparse-reward environment of Overcooked-AI, we applied reward shaping to augment raw rewards with an auxiliary-shaped signal.
 During reward shaping, each agent received rewards for performing specific actions aligned with the given recipe’s requirements: adding onions or tomatoes to the pot when an additional ingredient is needed, cooking the pot once it contains the correct quantity of ingredients, retrieving the soup, and finally serving it.
 The shaping reward $r_{s}$ was annealed over a predefined horizon ($2.5\times 10^{6}$) using a linear schedule.
-A.4 Experiment 1: Significance Tests
+Compute
+All experiments were run on A100 GPUs, totaling 1,569 GPU-hours across 17,813 runs.
+Experiment 1 used 93 GPU-hours (160 runs on MPE and SMAC).
+Experiment 2 used 120 GPU-hours (1,440 runs), Experiment 3 used 1,333 GPU-hours (16,000 runs), and Experiment 4 used 23 GPU-hours (280 runs).
+Appendix F Experiment 2: Results
+Number of Pots Mean SI Mean Reward
+1 0.676053 14.861111
+2 0.280481 18.867647
+Table 1: Average specialization (SI) with varying numbers of pots in Experiment 2.
+Spatial Layout Mean SI Mean Reward
+Open 0.3320525 18.7083330
+Horizontal Divider 0.332101 18.0416668
+Vertical Divider 0.318273 18.0000000
+Vertical Space 0.404964 18.0833335
+Long Divider 0.752011 12.2083334
+Two Dividers 0.2988275 17.6666666
+Table 2: Average specialization (SI) in various spatial layouts in Experiment 2.
+Appendix G Experiment 3: Significance Tests
 Number of Pots and Workstations.
@@
 Agents tended to specialize more with one pot ($M=0.77,SD=0.30$) than two ($M=0.57,SD=0.41$) but not when presented with more of other workstations (e.g., bowls, serving stations).
-A Welch’s t-test between layouts with one pot versus two revealed a significant difference in JSD, $t(2842.72)=14.98,p=7.23\times 10^{-49}$.
-Conversely, a one-way ANOVA showed a non-significant difference in JSD across the full range of possible workstation combinations ($F(5,3094)=0.82,p=0.54$).
+A Welch’s t-test between layouts with one pot versus two revealed a significant difference in SI, $t(2842.72)=14.98,p=7.23\times 10^{-49}$.
+Conversely, a one-way ANOVA showed a non-significant difference in SI across the full range of possible workstation combinations ($F(5,3094)=0.82,p=0.54$).
 Spatial Bottlenecks.
@@
 Larger centrality values represented more critical nodes, corresponding to greater spatial bottlenecks.
-A one-way ANOVA showed a significant difference in JSD across bins ($F(2,2097)=34.10,p=2.26\times 10^{-15}$).
-Pairwise comparisons using Tukey’s HSD showed that JSD was significantly higher with medium centrality than low ($p=3.18\times 10^{-7}$), high centrality than medium ($p=1.07\times 10^{-12}$), and high centrality than low ($p=3.82\times 10^{-20}$).
+A one-way ANOVA showed a significant difference in SI across bins ($F(2,2097)=34.10,p=2.26\times 10^{-15}$).
+Pairwise comparisons using Tukey’s HSD showed that SI was significantly higher with medium centrality than low ($p=3.18\times 10^{-7}$), high centrality than medium ($p=1.07\times 10^{-12}$), and high centrality than low ($p=3.82\times 10^{-20}$).
 Recipe Complexity.
-Finally, we found a significant effect of recipe complexity (the number of ingredients required to make a soup) on JSD.
+Finally, we found a significant effect of recipe complexity (the number of ingredients required to make a soup) on SI.
 Agents tended to specialize less when recipes had one ingredient ($M=0.54,SD=0.47$) than with two ($M=0.73,SD=0.35$) or three ($M=0.67,SD=0.32$).
-A one-way ANOVA showed a significant difference in JSD across recipes ($F(2,3097)=57.22,p=3.95\times 10^{-25}$).
-Pairwise comparisons using Tukey’s HSD revealed that JSD was significantly higher for recipes with two ingredients compared to one ($p=4.46\times 10^{-19}$) and recipes with three ingredients compared to one ($p=4.37\times 10^{-10}$), but significantly lower for recipes with three ingredients compared to two ($p=7.77\times 10^{-6}$).
-A.5 Experiment 1: Evaluation
+A one-way ANOVA showed a significant difference in SI across recipes ($F(2,3097)=57.22,p=3.95\times 10^{-25}$).
+Pairwise comparisons using Tukey’s HSD revealed that SI was significantly higher for recipes with two ingredients compared to one ($p=4.46\times 10^{-19}$) and recipes with three ingredients compared to one ($p=4.37\times 10^{-10}$), but significantly lower for recipes with three ingredients compared to two ($p=7.77\times 10^{-6}$).
+Appendix H Experiment 3: Evaluation
 For each of the 14 layouts, there were a total of 16 workstation combinations with three randomized positions each.
 We computed the predicted parallelizability, $S$, for each of these combinations and averaged over the three randomized locations to compute a measure of task parallelizability for each combination of layout, workstations, and recipe.
-We also fit a logistic regression model where the binary outcome variable was low versus high specialization (JSD $\geq 0.5$) and the predictor variable was task parallelizability.
-To address the class imbalance in the JSD distribution, we used class-weight balancing to automatically adjust class weights inversely proportional to their frequencies in the training data.
-Figure 6: Overcooked Spatial Layouts in Experiment 1.
-Figure 7: Effect of grid size on observed specialization (JSD) in Experiment 1.
+We also fit a logistic regression model where the binary outcome variable was low versus high specialization (SI $\geq 0.5$) and the predictor variable was task parallelizability.
+To address the class imbalance in the SI distribution, we used class-weight balancing to automatically adjust class weights inversely proportional to their frequencies in the training data.
+Figure 7: Overcooked Spatial Layouts in Experiment 3.
+Figure 8: Effect of grid size on observed specialization (SI) in Experiment 3.
 As the size of the grid increased, agends tended to specialize more (even when presented with the same number of workstations and bottlenecks).
-W
-Spatial Layout Mean JSD Mean Reward
+Spatial Layout Mean SI Mean Reward
 H_shape 0.666725 14.165217
@@
 two_islands 0.758321 13.133005
-Table 1: Average specialization (JSD) observed in different spatial layouts in Experiment 1.
-Workstations Mean JSD Std JSD
+Table 3: Average specialization (SI) observed in different spatial layouts in Experiment 3.
+Workstations Mean SI Std SI
 1 onions 0.69 0.38
@@
 2 bowls 0.67 0.39
-Table 2: Mean and standard deviation of observed specialization (JSD) for varying numbers of workstations in Experiment 1.
-Recipes Mean JSD Std JSD
+Table 4: Mean and standard deviation of observed specialization (SI) for varying numbers of workstations in Experiment 3.
+Recipes Mean SI Std SI
 1 onion 0.56 0.48
@@
 2 onions 1 tomato 0.77 0.29
-Table 3: Mean and standard deviation observed specialization (JSD) for different recipes in Experiment 1.
+Table 5: Mean and standard deviation observed specialization (SI) for different recipes in Experiment 3.
 Table 6: Correlations Across Spatial Layouts
@@
 H_shape -0.493 0.004
-L_shaped -0.699 $<0.001$
-U_shaped -0.679 $<0.002$
+L_shaped -0.699 <0.001<0.001
+U_shaped -0.679 <0.002<0.002
 big_room -0.568 0.001
@@
 counter_circuit -0.361 0.042
-cramped_room -0.619 $<0.001$
+cramped_room -0.619 <0.001<0.001
 gallery -0.524 0.002
@@
 Average Correlation -0.494
-Table 5: Correlations between predicted task parallelizability ($S$) and observed specialization (JSD) across varying spatial layouts in Experiment 1.
+Table 7: Correlations between predicted task parallelizability ($S$) and observed specialization (SI) across varying spatial layouts in Experiment 3.
 The significance of this relationship varied across layouts, especially those with different grid sizes, prompting the controlled design of Experiments 2 and 3.
-Figure 8: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific spatial layout.
-Figure 9: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution for a specific number of workstations.
-Figure 10: Specialization (JSD) distributions across all permutations in Experiment 1. Each subfigure represents the JSD distribution grouped by a specific recipe.
-Figure 11: Specialization (JSD) distributions across all random seeds in Experiment 1.
-Each subfigure represents the JSD distribution grouped by a specific layout.
+Figure 9: Specialization (SI) distributions across all permutations in Experiment 3. Each subfigure represents the SI distribution grouped by a specific spatial layout.
+Figure 10: Specialization (SI) distributions across all permutations in Experiment 3. Each subfigure represents the SI distribution for a specific number of workstations.
+Figure 11: Specialization (SI) distributions across all permutations in Experiment 3. Each subfigure represents the SI distribution grouped by a specific recipe.
+Figure 12: Specialization (SI) distributions across all random seeds in Experiment 3.
+Each subfigure represents the SI distribution grouped by a specific layout.
 The distributions of strategies appear mostly consistent across different random seeds.
 Five random seeds were used: 1, 2, 3, 5, and 7.
-Figure 12: Specialization (JSD) distributions across all random seeds in Experiment 1.
-Each subfigure represents the JSD distribution grouped by the number of workstations.
+Figure 13: Specialization (SI) distributions across all random seeds in Experiment 3.
+Each subfigure represents the SI distribution grouped by the number of workstations.
 The distributions appear consistent across the different random seeds.
 Five random seeds were used: 1, 2, 3, 5, and 7.
-Figure 13: Specialization (JSD) distributions across all random seeds in Experiment 1.
-Each subfigure represents the JSD distribution grouped by a specific recipe.
+Figure 14: Specialization (SI) distributions across all random seeds in Experiment 3.
+Each subfigure represents the SI distribution grouped by a specific recipe.
 The distributions appear consistent across the different random seeds.
 Five random seeds were used: 1, 2, 3, 5, and 7.
-A.6 Experiment 2: Results
-Number of Pots Mean JSD Mean Reward
-1 0.676053 14.861111
-2 0.280481 18.867647
-Table 6: Average specialization (JSD) with varying numbers of pots in Experiment 2.
-Spatial Layout Mean JSD Mean Reward
-Open 0.3320525 18.7083330
-Horizontal Divider 0.332101 18.0416668
-Vertical Divider 0.318273 18.0000000
-Vertical Space 0.404964 18.0833335
-Long Divider 0.752011 12.2083334
-Two Dividers 0.2988275 17.6666666
-Table 7: Average specialization (JSD) in various spatial layouts in Experiment 2.
-A.7 Experiment 3: Design
-Figure 14: Three layouts from Experiment 3 are shown: (a) the smallest layout, (b) the largest layout with increased exploration, and (c) the largest layout with consistent exploration.
-Across these configurations, the optimal policy remains constant.
-The plots depict how the JSD and reward vary with increasing layout size under two different conditions.
-In the case of the three-onion soup, the reward decreases as the layout size grows.
-The smallest layout (layout 1) corresponds to the minimum JSD and the maximum reward.
-In Experiment 3, we examined the influence of various divider configurations on agents specialization.
-Four skeleton layouts of identical size were tested: (1) a single vertical divider, (2) a double vertical divider, (3) a horizontal divider, and (4) no divider.
-The focal point was to investigate how these structural variations affect the convergence of agent strategies.
-To evaluate the impact of state space size on optimal policy convergence and in cases of suboptimal convergence, whether agents developed more specialist or generalist strategies - we conducted a controlled experiment comprised of 280 runs.
-The experiment systematically varied seven layout sizes, utilizing 10 random seeds per layout, two recipe configurations (1-onion soup and 3-onion soup), and two experimental conditions (constant vs.
-non-constant exploration).
-Critically, the optimal policy remained consistent across all runs; only empty space that the agents could not access was added to increase the size of the state space.
-Figure 14 illustrates three of these layouts, demonstrating how variations in layout size were implemented, while keeping the optimal policy consistent
-A.8 Experiment 3: Results
-For the three-onion soup task, the majority of run converged to a JSD value fo approximately 0.6 - achieved in $93.8\%$ of runs with a single divider, $87.5\%$ of runs with a double divider, $87.5\%$ in the no-divider room layout, and $93.8\%$ in the counter circuit layout.
-Notably, a JSD value of 0 corresponds to the optimal strategy - however, agents rarely reached this level of convergence.
-For the one-onion soup tasks, full specialization was observed in most runs across the four layouts; $100\%$ with a single divider, $81.3\%$ with a double divider, and $81.3\%$ with no-divider.
-In stark contrast, only $46.3\%$ of runs in the counter circuit layout converged to a specialist strategy ($JSD>0.5$), while the remaining $53.4\%$ of runs resulted in generalist behavior.
-The differences in emergent strategies between layouts is likely due to the distinct effects of the dividers on agent behavior.
-In layouts with vertical dividers, agents are more inclined to specialize, as the dividers channel them toward the same region where they cannot work in parallel.
-Conversely, the counter circuit layout encourages generalist strategies by focusing them to either occupy separate regions or follow each other, reducing the likelihood of specialization.
-The effect of layouts size varied between the two recipes.
-For the simpler 1-onion soup recipe, there was no significant relationship between layout size and either reward JSD.
-Under non-constant exploration, JSD increased significantly with layout size ($r=0.7924,p=0.0336$) indicating greater policy specialisation in large layouts, but reward showed no reliable trend ($r=-0.2512,p=0.5869$).
-Pairwise T-tests confirmed no significant difference between the smallest and largest environments for JSD ($p=0.755$ or reward $p=0.210$.
-Under constant exploration, the 1-onion soup was unaffected by layout size, with no substantial correlation for JSD ($r=-0.5850,p=0.1677$) or reward ($r=0.0609,p=0.8969$).
-In contrast, the 3-onion soup recipe was sensitive to increasing layout size.
-Under non-constant exploration, as layout size increased, JSD rose significantly ($r=0.7549,p=0.0498$) while reward decreased ($r=-0.8556,p=0.0140$), indicating that agents struggle to perform well in large environments and develop more specialised policies.
-Pairwise T-tests further outlined significant differences between the smallest and largest layouts for JSD ($p=0.008$ and reward $p<10^{-4}$.
-Even under constant exploration, the 3-onion soup recipe showed a decline in reward as layout size increased ($r=-0.4308,p=0.0020$), although JSD was not significantly affected.
-These findings demonstrate that increasing both layout size and the number of subtasks, thus expanding the exploration space, drives agents toward suboptimal and overly specialized policies.
-Even when exploration remains constant, the addition of extraneous state dimensions in larger layouts prevents the RL policy’s ability to converge effectively to the optimal solution.
-This underscores a key challenge for future work: MARL policies struggle to consistently identify optimal generalist solutions in larger layouts, even when task structure, workstations, and exploration dynamics remain unchanged.
-Layout Experiment Recipe Mean JSD Std JSD Count Mean Reward Std Reward Reward Count
+Appendix I Experiment 4: Results
+Layout Experiment Recipe Mean SI Std SI Count Mean Reward Std Reward Reward Count
 1 A 1-Onion 0.750219 0.397342 10 24.8 6.014797 10
@@
 7 B 3-Onion 0.713627 0.403798 5 6.5 3.439961 10
-Table 8: Detailed results from Experiment 3.
-Each row represents a unique combination of spatial layout, experiment, and recipe, showing the mean, standard deviation (std), and count (number of runs) for both specialization (JSD) and reward metrics.
+Table 8: Detailed results from Experiment 4.
+Each row represents a unique combination of spatial layout, experiment, and recipe, showing the mean, standard deviation (std), and count (number of runs) for both specialization (SI) and reward metrics.
 Experiment A corresponds to expanding exploration, while Experiment B denotes constant exploration.
-The JSD count is less than 10 for certain permutations because the policy failed to converge, resulting in the agents taking no meaningful actions.
+The SI count is less than 10 for certain permutations because the policy failed to converge, resulting in the agents taking no meaningful actions.
 Recipe Exploration Metric Corr. p-value
-1-Onion Expanding JSD 0.7924 0.0336
+1-Onion Expanding SI 0.7924 0.0336
 Reward -0.2512 0.5869
-Constant JSD -0.5850 0.1677
+Constant SI -0.5850 0.1677
 Reward 0.0609 0.8969
-3-Onion Expanding JSD 0.7549 0.0498
+3-Onion Expanding SI 0.7549 0.0498
 Reward -0.8556 0.0140
-Constant JSD -0.0423 0.9283
+Constant SI -0.0423 0.9283
 Reward -0.6947 0.0832
-Table 9: Correlations from Experiment 3.
-Correlations between specialization (JSD) and reward for one-onion and three-onion Soup recipes under expanding and constant exploration conditions.
-The table reports correlations and p-values ($p$) for JSD and reward in relation to layout size.
+Table 9: Correlations from Experiment 4.
+Correlations between specialization (SI) and reward for one-onion and three-onion Soup recipes under expanding and constant exploration conditions.
+The table reports correlations and p-values ($p$) for SI and reward in relation to layout size.
 Condition Metric t-value p-value
 Combined Recipes
-layout_1 vs layout_7 JSD -2.229 0.029
+layout_1 vs layout_7 SI -2.229 0.029
 layout_0 vs layout_7 Reward 0.506 0.614
 1-Onion Soup Recipe
-layout_1 vs layout_7 JSD -0.315 0.755
+layout_1 vs layout_7 SI -0.315 0.755
 layout_1 vs layout_7 Reward -1.278 0.210
 3-Onion Soup Recipe
-layout_1 vs layout_7 JSD -2.820 0.008
+layout_1 vs layout_7 SI -2.820 0.008
 layout_1 vs layout_7 Reward 5.812 0.000
-Table 10: Results from Experiment 3. Pairwise T-tests comparing the smallest (Layout 1) layout with the largest layout (Layout 67).
+Table 10: Results from Experiment 4. Pairwise T-tests comparing the smallest (Layout 1) layout with the largest layout (Layout 67).